### v1

# LLaVA-MoD: Making LLaVA Tiny via MoE-Knowledge Distillation

Abstract
We introduce LLaVA-MoD, a novel framework designed to enable the efficient training of small-scale Multimodal Language Models (s-MLLM) by distilling knowledge from large-scale MLLM (l-MLLM). Our approach tackles two fundamental challenges in MLLM distillation. First, we optimize the network structure of s-MLLM by integrating a sparse Mixture of Experts (MoE) architecture into the language model, striking a balance between computational efficiency and model expressiveness. Second, we propose a progressive knowledge transfer strategy to ensure comprehensive knowledge migration. This strategy begins with mimic distillation, where we minimize the Kullback-Leibler (KL) divergence between output distributions to enable the student model to emulate the teacher network’s understanding. Following this, we introduce preference distillation via Direct Preference Optimization (DPO), where the key lies in treating l-MLLM as the reference model. During this phase, the s-MLLM’s ability to discriminate between superior and inferior examples is significantly enhanced beyond l-MLLM, leading to a better student that surpasses its teacher, particularly in hallucination benchmarks. Extensive experiments demonstrate that LLaVA-MoD outperforms existing models across various multimodal benchmarks while maintaining a minimal number of activated parameters and low computational costs. Remarkably, LLaVA-MoD, with only 2B activated parameters, surpasses Qwen-VL-Chat-7B by an average of 8.8% across benchmarks, using merely $0.3\%$ of the training data and 23% trainable parameters. These results underscore LLaVA-MoD’s ability to effectively distill comprehensive knowledge from its teacher model, paving the way for the development of more efficient MLLMs. The code will be available on https://github.com/shufangxun/LLaVA-MoD
Figure 1: Comparisons of training cost and performance. LLaVA-MoD achieves competitive performance with advanced MLLMs using significantly lower training costs while outperforming current small-scale MLLMs by a large margin.
1 Introduction
Leveraging the advanced instruction-following and reasoning capabilities of Large Language Models (LLMs) (Achiam et al., 2023; Bai et al., 2023a; Jiang et al., 2024; Team et al., 2023; Touvron et al., 2023a; b; Dubey et al., 2024), Multimodal Large Language Models (MLLMs) (Bai et al., 2023b; Liu et al., 2024; Lin et al., 2024b; Li et al., 2023b; Chen et al., 2023b; Shu et al., 2023; Lu et al., 2024), typically integrate a visual encoder (Caron et al., 2021; Radford et al., 2021; Liu et al., 2022) alongside an LLM to achieve promising results across various vision and cross-modal tasks, such as image captioning and visual question-answering. MLLMs are characterized by their large model size and extensive training data, which contribute to their superior performance but also necessitate significant computation resources. For instance, the largest version of LLaVA-NeXT (Li et al., 2024a) employs Qwen-1.5-110B (Yang et al., 2024) as the language model and requires 18 hours of training with 128 H800 GPUs. Furthermore, MLLMs face deployment challenges due to the high memory and computational demands. For instance, models with substantial parameters often require advanced hardware and exhibit slow inference speeds, which significantly hinder their real-world applications such as on mobile devices. Thus, developing MLLMs that balance performance and efficiency has become a crucial research focus.
Previous works on small-scale Multimodal Large Language Models (s-MLLM) (Zhou et al., 2024a; Yuan et al., 2023; Shao et al., 2024; He et al., 2024; Chu et al., 2023; 2024; Yao et al., 2024) have mainly concentrated on devising data collection (Schuhmann et al., 2021; 2022; Zhu et al., 2024; Awadalla et al., 2024; Gadre et al., 2024; Chen et al., 2023a) and filtering protocols (Xu et al., 2023; Fan et al., 2024; Zhang et al., 2023; Li et al., 2024b; Nguyen et al., 2024) to ensure high-quality training data. While such high-quality data aid s-MLLM in narrowing the performance gap with their larger counterparts, their effectiveness is still limited by model capacity and the availability of high-quality data. With the advent of open-source MLLMs, an intuitive strategy inspired by conventional deep learning tasks involves utilizing large-scale MLLMs (l-MLLM) as the teacher to improve the training of s-MLLM via Knowledge Distillation (KD) (Hinton et al., 2015). KD facilitates stable and efficient training of smaller models by aligning the output distributions of the student model with those of the teacher model. However, the application of KD as a model reduction technique has yet to be extensively explored in MLLMs. To develop an effective framework for MLLM distillation, we consider two primary challenges. The first challenge lies in designing a lightweight architecture for student MLLMs that retains strong learning and expressive capabilities, enabling it to absorb the complex knowledge embedded within teacher MLLMs effectively. The second challenge is to efficiently and comprehensively transfer this knowledge from the teacher MLLMs to the student MLLMs.
To address these two challenges, we present LLaVA-MoD [^1] MoD denotes Mixture-of-Expert Knowledge Distillation for MLLM, an effective framework for training s-MLLM by mimicking the behavior of the l-MLLM through Mixture-of-Expert (MoE) Knowledge Distillation. For the first challenge—designing a lightweight s-MLLM structure—a straightforward approach might involve reducing the scale of the base language model in l-MLLM to create a smaller network. However, this direct reduction in network parameters can significantly compromise the model’s expressive capability, making it less effective in modeling complex multimodal understanding tasks. Drawing inspiration from the recent success of sparse MoE (Lin et al., 2024a; Dai et al., 2024; Jiang et al., 2024; Shen et al., 2023) in sequential modeling, we incorporate a sparse MoE structure into the dense s-MLLM. This approach seeks to balance scale reduction while preserving the model’s ability to capture and represent intricate multimodal information during distillation. Specifically, we enhance the s-MLLM by equipping it with several feedforward networks (FFNs) and linear gates within the LLM. Each FFN acts as an expert, capturing fine-grained knowledge from the l-MLLM, while the gates select the top-$k$ experts to facilitate optimal knowledge transfer pathways.
To address the second challenge, we propose a progressive distillation strategy for effective knowledge transfer. The process begins by aligning the vision encoder with the LLM using a learnable adapter, thereby initializing a dense student MLLM. Following this, we employ two consecutive distillation stages, where the student model evolves from mimicking and approximating the teacher MLLM to ultimately surpassing it: Mimic Distillation. This stage is divided into two steps, i.e., dense-to-dense and dense-to-sparse distillation. In the dense-to-dense distillation stage, we employ standard knowledge distillation loss to align the output logits distribution between the initialized dense s-MLLM and the teacher model, using general captioning and conversation datasets. Next, we transition to the dense-to-sparse distillation where we transform the dense student MLLM into a sparse one by integrating MoE, and then distill knowledge from the teacher MLLM into this sparse student model using a broad range of tasks and datasets. Preference Distillation. In this stage, the teacher model provides knowledge regarding what constitutes “good” and “bad” samples, establishing a foundational reference for the student model. The student MLLM leverages this knowledge to adjust its probability distribution, ensuring that good samples have a higher probability than those from the teacher model, while bad samples are assigned a lower probability. This process enhances the student model’s ability to mitigate hallucinations by improving its judgment capabilities beyond those of the teacher model.
As illustrated in Figure 1, LLaVA-MoD exhibits impressive performance across various multimodal benchmarks while maintaining minimal activated parameters and low computational resources. For example, LLaVA-MoD-2B exceeds Qwen-VL-Chat-7B by an average of $8.8\%$ on these benchmarks, utilizing just $0.3\%$ of the training data and $23\%$ trainable parameters. Furthermore, it matches the performance of RLHF-based methods with 7B and 13B parameters on several hallucination benchmarks. Specifically, LLaVA-MoD-2B surpasses RLHF-V (Yu et al., 2024a) by $8.2\%$ in response-level hallucination rate and by $21.3\%$ in mention-level hallucination rate on the Object HalBench. The impressive results demonstrate the effectiveness of our proposed MoD framework in transferring knowledge from l-MLLM to s-MLLM.
2 Related Work
Multimodal Large Language Models. The emergence of large language models (LLMs) has greatly advanced the field of natural language processing. Connecting visual information to LLM to enhance their understanding of multimodal input is crucial for promoting the unified comprehension of vision and language. CLIP (Radford et al., 2021) was the first to align visual and textual information into a unified embedding space through contrastive learning objectives. BLIP-2 (Li et al., 2023b) utilizes various pre-training tasks to train separate visual and text streams, and adds additional intermediary structures to adapt visual features to LLM. Flamingo (Alayrac et al., 2022) incorporates additional cross-attention modules into LLM to handle arbitrarily interleaved sequences of images and texts. Recently, models like LLaVA (Liu et al., 2024) and MiniGPT-4 (Zhu et al., 2023) use linear projection layers to map image features to the textual space and enhance the models’ instruction-following ability through visual instruction tuning. Moreover, some works focus on a stronger vision encoder. Intern-VL (Chen et al., 2023b) leverages a larger-scale vision encoder to bridge the gap with a large-scale language model, and Mini-Gemini (Li et al., 2024c) introduces additional vision encoders for high-resolution refinement to enhance visual tokens. Qwen-VL (Bai et al., 2023b), VisionLLM (Wang et al., 2023c), and others aim to imbue models with stronger fine-grained visual understanding capabilities, such as grounding and region understanding. Unlike these approaches, our method does not emphasize enlarging the model to boost multimodal capabilities. Instead, we adopt a distillation approach to compress the capabilities of larger models into a smaller, sparse MoE architecture, enhancing computational and storage efficiency.
Knowledge Distillation. Large language models possess powerful capabilities, but their enormous size and high inference costs limit their application in low-resource environments. Knowledge distillation (Hinton et al., 2015) uses a large model as teacher to transfer its advanced knowledge to a smaller student model, which plays a critical role in compressing model size and enables smaller models to self-improve. Some works adopt additional designs to make the distillation process more suitable for LLMs. MiniLLM (Gu et al., 2023) minimizes the reverse KLD (Kullback–Leibler divergence) to prevent students from overestimating low-probability regions in the teacher distribution, while GKD (Agarwal et al., 2023) introduces generalized knowledge distillation and promotes the integration of distillation with RLHF (Reinforcement Learning from Human Feedback). Additionally, Huang et al. (Huang et al., 2022) introduce context distillation, and some works (Mukherjee et al., 2023; Li et al., 2022; Ho et al., 2023) adopt CoT (Chain-of-Thought) distillation to enhance specific skills of small models. Our approach innovatively distills the knowledge of MLLM into a smaller sparse MoE (Mixture-of-Experts) architecture, significantly enhancing the multimodal processing capabilities of small models at a minimal cost.
Mixture-of-Experts The mixture of experts (MoE) architecture is initially introduced by Jacobs et al. (1991) to enhance model performance across various samples through the utilization of independent experts. Recently, with the rise of transformer architectures, numerous studies have extended transformer layers to incorporate MoE. In transformer-based MoE architecture, feed-forward neural network (FFN) layers are often replaced with sparsely activated experts, employing a trainable Top-k gating strategy (Lepikhin et al., 2020; Fedus et al., 2022), thereby effectively increasing model capacity while maintaining a low computational overhead. Additionally, to further mitigate training costs, the sparse upcycling method (Komatsuzaki et al., 2022) is adopted, which initializes expert parameters using those from a well-trained dense model. Currently, MoE architectures have found widespread application not only in pre-trained language models (Jiang et al., 2024; Dai et al., 2024) but also in vision models and vision-language models (Lin et al., 2024a; Shen et al., 2023). Our approach integrates MoE with knowledge distillation techniques to provide stronger signals for sparse training, which remarkably decrease the training cost associated with sparse models.
3 Method
Figure 2: Progressive Distillation of LLaVA-MoD. (a). Mimic Distillation: Aligning the student’s response probabilities $p_{S}$ with those of the teacher $p_{T}$, via the Kullback–Leibler (KL) Loss. (b). Preference Distillation: Increasing the student’s positive response probabilities $p^{+}_{S}$ to surpass those of the teacher $p^{+}_{T}$, while decreasing the student’s negative response probabilities $p^{-}_{S}$ to fall below those of the teacher $p^{-}_{T}$, via the Preference-Optimization (PO) Loss.
We introduce LLaVA-MoD, a novel framework for building efficient MLLM using mixture-of-experts (MoE) knowledge distillation. Our framework consists of two main components: (a). Architecture Design of s-MLLM: As shown in Figure 3, we design the s-MLLM with a sparse MoE framework, enhancing the ability to acquire specialized expert knowledge. (b). Distillation Mechanism: We design a progressive distillation mechanism as shown in Figure 2 to transfer knowledge from the l-MLLM to the sparse s-MLLM. This process involves two stages: mimic distillation followed by preference distillation. We will introduce LLaVA-MoD in detail from these two components.
3.1 Architecture Design of Sparse s-MLLM
In this section, we describe the architecture design of our sparse s-MLLM, which serves as the student model in the distillation process.
s-MLLM Definition.
As illustrated in Figure 3, the basic architecture of s-MLLM consists of three primary components: a vision encoder, a large language model (LLM), and a vision-language (VL) adaptor. Given a multimodal instruction conversation $(x,y)$, we define our s-MLLM to process response $y$ as follows:
$y=\text{LLM}_{\phi}(\text{Proj}_{\omega}(\text{ViT}_{\chi}(x_{v})),x_{i}),$ (1)
where $x_{v}$ is the input image, and $x_{i}$ is the text instruction. The input image is resized to 336 $\times$ 336 and patched into 576 image tokens, each of size 14 $\times$ 14. $\text{ViT}_{\chi}$ is the CLIP vision encoder with parameters $\chi$ which first extracts image features from $x_{v}$. $\text{Proj}_{\omega}$ is the vision-language adaptor with parameters $\omega$, serving as the vision tokenizer to align the image features with the word embedding space. $\text{LLM}_{\phi}$ is the large language model with parameters $\phi$, which produces the response $y$ based on the multimodal tokens of $x=[x_{v},x_{i}]$.
Figure 3: Dense to Sparse Architecture of LLaVA-MoD. The architecture of LLaVA-MoD consists of three components: VL Adaptor, Vision Encoder, and LLM. The VL Adaptor and Vision Encoder remain unchanged, while the LLM is upcycled from dense to sparse.
Mixture-of-Experts.
The principle of building our s-MLLM is downsizing the LLM while leaving the vision encoder and vision-language adaptor unmodified. To achieve this downsizing goal, we sparsify the dense s-MLLM by incorporating an MoE architecture. Specifically, Figure 3 illustrates the process, where we apply the sparse upcycling technique (Komatsuzaki et al., 2022) to replicate $N$ feedforward networks (FFNs) as the expert modules. Additionally, we introduce a linear layer as the router, which dynamically activates the appropriate experts by predicting the probability of expert assignment. Given each token $x$ in the sequence, we first compute the routing value of $N$ experts:
$r=\text{Softmax}(x\cdot W_{r}),$ (2)
where $W_{r}$ denotes the weight matrix of the router and each element $r_{i}$ in $r$ represents the probability of activating the $i$-th expert. After that, we apply the Top-$k$ strategy to determine the activated experts with the highest $k$ routing values:
$\tilde{r}=\text{Top-}k({r})=\begin{cases}r_{i},&\text{if }i\in k\\ 0,&\text{otherwise}\\ \end{cases}$ (3)
where the routing values for the inactivated experts are set to zero, effectively excluding them from contributing to the final output. The output $y$ is calculated by aggregating the contributions of the activated experts, weighted by their corresponding routing values:
$y=\sum_{i=1}^{N}\tilde{r}\cdot E_{i}(x).$ (4)
3.2 Progressive Distillation
Our progressive distillation consists of two distinct stages, i.e., mimic distillation (Figure 2 (a)) and preference distillation (Figure 2 (b)). In the mimic distillation stage, the student MLLM $\pi_{\text{S}}$ imitates the general and specific knowledge from the teacher MLLM $\pi_{\text{T}}$. In the preference distillation stage, the student MLLM gains the preference knowledge of the teacher MLLM to further refine its output and reduce hallucinations. We use the pre-trained l-MLLM as the teacher MLLM, which is frozen during the whole distillation process. Both teacher and student MLLMs are in the same LLM family, ensuring consistent vocabulary space, which is essential for the student MLLM to imitate the teacher MLLM accurately. For simplicity, we denote the student as $\pi_{\text{S}}$ and the teacher as $\pi_{\text{T}}$.
Initialization.
Before distillation, we first align the vision encoder with the LLM via a learnable adapter, aiming to obtain a well-initialized dense student MLLM. The $\text{LLM}_{\phi}$ and $\text{ViT}_{\chi}$ are maintained frozen since their pre-trained parameters have already captured rich visual and language knowledge. Only $\text{Proj}_{\omega}$ is optimized to bridge the gap between the vision and language domain. For the initialization, we utilize common image-caption pairs from a widely used and curated dataset, which covers a diverse range of subjects and visual entities. The training objective is to minimize the cross-entropy of the generated tokens. The objective function is defined as:
$\mathcal{L}_{\text{Init}}(\pi_{\text{S}})=-\mathbb{E}_{(y_{k}\mid y_{<k},x)\sim\pi_{\text{S}}}\left[\log\pi_{\text{S}}(y_{k}\mid y_{<k},x)\right],$ (5)
where $\pi_{\text{S}}(y_{k}\mid y_{<k},x)$ represents the probability of the predicted token $y_{k}$ conditioned on $x$ and the previous tokens $y_{<k}=(y_{1},y_{2},\dots,y_{k-1})$.
Mimic Distillation.
We decompose the comprehensive knowledge of the teacher MLLM into general and expertise aspects and undertake dense-to-dense distillation and dense-to-sparse distillation respectively to transfer this knowledge to the student MLLM.
a). Dense-to-Dense Distillation. In this stage, we aim to replicate the general knowledge of the teacher MLLM. Acquiring general knowledge first is crucial, as it provides a broad foundation and common understanding across various domains, enabling the student MLLM to develop a basic framework applicable in multiple scenarios. This foundation supports a more comprehensive and flexible understanding before progressing to specialized tasks. To achieve this, we maintain $\text{ViT}_{\chi}$ frozen, and jointly optimize $\text{LLM}_{\phi}$ and $\text{Proj}_{\omega}$, with trainable parameters $\theta=\{\omega,\phi\}$. We leverage common image-caption pairs and conversation datasets. The training objective is to minimize the Kullback-Leibler divergence (KLD) between the output logits of the dense student MLLM and the teacher MLLM. The objective function is defined as:
$\mathcal{L}_{\text{D2D}}(\pi_{\text{S}};\pi_{\text{T}})=-\mathbb{E}_{(x,y_{k})\sim\pi_{\text{T}}}\left[\log\frac{\pi_{\text{T}}(y_{k}\mid y_{<k},x)}{\pi_{\text{S}}(y_{k}\mid y_{<k},x)}\right],$ (6)
where $V$ denotes the vocabulary, while $\pi_{\text{S}}(y_{k}\mid y_{<k},x)$ and $\pi_{\text{T}}(y_{k}\mid y_{<k},x)$ denote the probability of the predicted tokens for student and teacher models, respectively.
b). Dense-to-Sparse Distillation. In this stage, our focus shifts to transferring the specialized expertise of the teacher MLLM, allowing the student MLLM to obtain advanced capabilities and achieve superior performance in diverse tasks. However, directly learning this knowledge in its dense form could lead to inefficiencies and difficulty. Therefore, we sparsify the dense student MLLM by introducing multiple experts. As detailed in Section 3.1, we replicate $N$ feedforward networks (FFNs) within $\text{LLM}_{\phi}$ and incorporate an MLP layer as a router, which jointly constitutes the experts with parameters $\phi_{e}$. This sparse architecture allows the student MLLM to selectively activate the most relevant experts for different tasks and inputs, offering significant advantages in emulating the teacher’s specialized knowledge. For training, we leverage multi-task data, updating only the experts and the adaptor. We employ a Top-$k$ routing strategy to select the experts. The trainable parameters are $\theta=\{\omega,\phi_{e}\}$. Similar to the previous stage, we also adopt the KLD as the training objective. Additionally, we include the standard next-token prediction objective, i.e., minimizing the student-generated tokens’ cross-entropy loss, to inject supervision from ground truth data which can reduce the existing bias in the teacher model. The final objective function is defined as:
$\mathcal{L}_{\text{D2S}}(\pi_{\text{S}};\pi_{\text{T}})=-\mathbb{E}_{(x,y_{k})\sim\pi_{\text{S}}}\left[\log\pi_{\text{S}}(y_{k}\mid y_{<k},x)\right]-\mathbb{E}_{(x,y_{k})\sim\pi_{\text{T}}}\left[\log\frac{\pi_{\text{T}}(y_{k}\mid y_{<k},x)}{\pi_{\text{S}}(y_{k}\mid y_{<k},x)}\right].$ (7)
Preference Distillation.
In this stage, our goal is to distill the preference knowledge from the teacher MLLM to guide the student MLLM towards generating not only accurate but also reasonable responses, which is crucial in reducing hallucinations. For the training process, we make effective use of preference data, which comprises meticulously paired positive responses $y^{+}$ and negative responses $y^{-}$ for the identical prompt $x$. Our preference distillation strategy is inspired by recent advancements in Direct Preference Optimization (DPO) (Rafailov et al., 2024), which bypasses the need for training a reward model by directly training on an offline preference dataset. Our key insight is to treat the teacher MLLM as the reference model during distillation. The teacher model plays a crucial role by providing insights into what constitutes “good” and “bad”, thereby establishing a fundamental reference for the student model.
Specifically, the training objective is to optimize the student model to assign higher probabilities to positive responses than those of the teacher model while assigning lower probabilities to negative responses than those of the teacher model. This training process involves two key optimization aspects: First, the student model seeks to align with the teacher model in distinguishing positive from negative responses. Second, the student model aims to surpass the teacher model by assigning higher probabilities to positive responses and lower probabilities to negative responses. We only train the mixture of experts and the VL adaptor in the student model and employ a Top-$k$ routing strategy to select the experts. The trainable parameters are $\theta=\{\omega,\phi_{e}\}$. The object function is defined as:
$\mathcal{L}_{\text{PD}}(\pi_{\text{S}};\pi_{\text{T}})=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\text{S}}(y^{+}\mid x)}{\pi_{\text{T}}(y^{+}\mid x)}-\beta\log\frac{\pi_{\text{S}}(y^{-}\mid x)}{\pi_{\text{T}}(y^{-}\mid x)}\right)\right],$ (8)
where $\pi_{\text{S}}(y^{+}\mid x)$ and $\pi_{\text{S}}(y^{-}\mid x)$ denote the probabilities of positive and negative responses for the student model, respectively, and $\pi_{\text{T}}(y^{+}\mid x)$ and $\pi_{\text{T}}(y^{-}\mid x)$ denote the same for the teacher model.
4 Experiments
4.1 Experimental Settings
Implementation Details.
We employ the well-established ”ViT-MLP-LLM” architecture to demonstrate the effectiveness of LLaVA-MoD. Specifically, a pre-trained CLIP-ViT-L/14 is utilized as the vision encoder and a 2-layer MLP is utilized as the adaptor for both the student and the teacher models. Both teacher and student models leverage the Qwen-1.5 and Qwen-2 families to construct their base model. Specifically, the teacher model employs a 7B parameter configuration, while the student model is constructed with 1.8B and 0.5B parameter sizes. The performance of teacher MLLM on common multimodal benchmarks is presented in Table 1.
Table 1: Performance of teacher MLLM on multimodal benchmarks.
[…]
Qwen-1.5-7B CLIP-L 62.3 40.7 70.9 55.3 72.1 68.5 64.9 62.1
Qwen-2-7B CLIP-L 62.5 37.9 73.2 57.2 78.0 71.6 71.1 64.5
As introduced in Section 3.2, the training process comprises mimic distillation and preference distillation, each with specific settings. At initialization, we first freeze the vision encoder and LLM while optimizing the VL adaptor to align image tokens with the word embedding space. This stage employs a cross-entropy loss with a batch size of 512 and a learning rate of 1e-4. In the mimic distillation stage, the vision encoder remains frozen, while the LLM and VL adaptor are co-optimized to distill general knowledge from the teacher MLLM in a dense-to-dense fashion. Then, the FFN in the LLM is first transformed into a sparse architecture with a mixture of FFN experts co-optimized with the VL adaptor to distill specialized knowledge from the teacher MLLM in a dense-to-sparse fashion. This stage employs a KL-divergence loss, with cross-entropy loss added for dense-to-sparse distillation. The batch size is 256, and the learning rate is adjusted to 2e-5. In the preference distillation stage, the model inherits from the mimic distillation. The vision encoder remains frozen, and a mixture of FFN experts is co-optimized with the VL adaptor to distill preference knowledge from the teacher MLLM. This stage employs a DPO loss to optimize the student’s probability for positive responses to be greater than that of the teacher and the student’s probability for negative responses to be lower than that of the teacher. The batch size is 256, and the learning rate is adjusted to 2e-6. Throughout all stages, we employ the adam optimizer (Diederik, 2014) and train on 16 NVIDIA A100 GPUs for one epoch each, totaling approximately 960 GPU hours. The detailed hyperparameters of each training stage are illustrated in Appendix  11.
Training Datasets.
The training data consists of 5M samples from the open-source datasets, with each training stage utilizing a distinct dataset. During Initialization, 0.6M general captioning samples from LLaVA-1.5-pretrain dataset (Liu et al., 2023b) are used to bridge the gap between visual and language modalities. In mimic distillation, 2.4M general captioning and conversation samples are used to distill general knowledge from the teacher MLLM and 1.4M multi-tasks data, including VQA, documents, science, and OCR, are used to distill specialized knowledge from the teacher MLLM. For preference distillation, 0.8M preference data samples are used to transfer preference knowledge from the teacher. The detailed dataset of each training stage is illustrated in Appendix 10.
Evaluation Benchmarks.
[…]
4.2 Main Results
In this section, we conduct experiments of LLaVA-MoD to highlight its advantages in two aspects: performance and efficiency. For performance, we evaluate comprehension-oriented benchmarks (Table 2) and hallucination-oriented benchmarks (Table 3). For efficiency, we present a comparison in terms of data samples and model size.
[…]
BLIP-2 Vicuna-13B 129M $\geq$7B 41.0 19.6 61.0 42.5 64.7 - - -
VILA-7B LLaMA-7B 50M 62.3 57.8 68.2 64.4 76.7 68.9 61.7 65.7
CogVLM Vicuna-7B 1500M 64.9 - 65.6 78.2 71.8 63.7 53.8 -
InstructBLIP Vicuna-13B 130M 49.5 33.4 63.1 50.7 60.6 - - -
IDEFICS-80B LLaMA-65B 6.8M 38.4 35.5 59.9 25.9 64.2 48.2 25.2 42.5
Qwen-VL-Chat Qwen-7B 1450M 57.5 38.9 68.2 61.5 74.4 60.6 56.7 56.7
SPHINX-Intern2 InternLM2-7B 15M 56.2 49.6 70.4 58.1 63.2 57.9 - -
Deepseek-VL-7B DLLM-7B 2000M 61.3 49.9 74.0 64.7 73.4 74.1 72.8 67.2
LLaVA-1.5-7B Vicuna-1.5-7B 1.2M 62.0 50.0 66.8 58.2 75.5 64.3 58.3 62.1
LLaVA-NeXT Vicuna-1.5-13B 1.3M 65.4 60.5 73.6 67.1 78.7 70.0 64.4 68.5
Imp-3B Phi-2-2.7B 1.6M $\sim$3B 63.5 54.1 72.8 59.8 72.3 72.9 46.7 63.2
Bunny-3B Phi-2-2.7B 2.7M 62.5 43.8 70.9 56.7 74.4 68.6 37.2 59.2
VILA-3B LLaMA-2.7B 51M 61.5 53.5 69.0 60.4 72.1 63.4 52.7 61.8
LLaVA-Phi Phi-2-2.7B 1.3M - 35.9 68.4 48.6 66.7 59.8 - -
TinyLLaVA Phi-2-2.7B 3.2M 61.0 - 70.1 53.5 71.6 68.3 - -
TinyGPT-V Phi-2-2.7B 24M 33.6 33.4 - - - - - -
MobileVLM MLLaMA-2.7B 1.3M 59.0 - 61.0 47.5 64.4 59.6 - -
MobileVLMv2 MLLaMA-2.7B 3.6M 61.1 - 70.0 57.5 72.0 63.2 - -
MoE-LLaVA-3B Phi-2-2.7B 2.2M 61.4 43.9 68.5 51.4 71.1 65.2 41.8 57.6
MiniCPM-V MiniCPM-2.4B 570M 51.5 50.5 74.4 56.6 68.9 64.0 62.7 61.2
MiniCPM-V-2 MiniCPM-2.4B 570M 52.1 60.2 76.3 73.2 70.5 68.5 67.2 66.9
Imp-2B Qwen-1.5-1.8B 1.6M $\sim$2B 61.9 39.6 66.1 54.5 65.2 63.8 61.2 58.9
Bunny-2B Qwen-1.5-1.8B 2.7M 59.6 34.2 64.6 53.2 65.0 59.1 58.5 56.3
Mini-Gemini-2B Gemma-2B 2.7M 60.7 41.5 63.1 56.2 67.0 59.8 51.3 57.1
MoE-LLaVA-2B Qwen-1.5-1.8B 2.2M 61.5 32.6 63.1 48.0 64.6 59.7 57.3 55.3
DeepSeek-VL-1.3B DLLM-1.3B 2000M 59.3 36.8 64.2 58.4 65.3 64.6 61.0 58.5
LLaVA-MoD-2B Qwen-1.5-1.8B 5M 58.7 39.2 68.0 58.5 66.7 66.3 61.9 59.9
LLaVA-MoD-2B Qwen-2-1.5B 5M 58.8 40.4 69.2 59.9 69.2 68.9 64.4 61.6
SPHINX-Tiny TLLaMA-1.1B 15M $\sim$1B 58.0 49.2 21.5 57.8 63.1 56.6 37.8 49.2
LLaVA-MoD-1B Qwen-1.5-0.5B 5M $\sim$1B 56.2 31.6 62.8 53.9 65.3 58.8 50.4 54.1
LLaVA-MoD-1B Qwen-2-0.5B 5M 56.6 35.1 61.1 57.1 67.0 58.7 54.1 55.7
4.2.1 Comprehension-Oriented Benchmarks
As indicated in Table 2, LLaVA-MoD demonstrates remarkable performance on comprehension-oriented benchmarks. It achieves SOTA average results among the models of 2B size and 1B size. Specifically, The 2B-sized LLaVA-MoD surpasses Mini-Gemini-2B (Li et al., 2024c) by 8.1%, while using a lower image resolution (336 vs. 768). The 1B-sized LLaVA-MoD surpasses SPHINX-Tiny (Gao et al., 2024) by 13.2%, using fewer data samples (5M vs. 15M). Furthermore, LLaVA-MoD-2B matches and even surpasses the performance of large-scale MLLMs. Specifically, The 2B-sized LLaVA-MoD surpasses Qwen-VL-Chat-7B (Bai et al., 2023b) by 8.8% and matches the performance of VILA-3B (Lin et al., 2024b) and MiniCPM-V (Yao et al., 2024). These results highlight that our approach effectively and efficiently training small-scale MLLMs by distilling the sparse MoE architecture.
Table 3: Comparison with state-of-the-art MLLMs on the hallucination benchmarks. We compare LLaVA-MoD with SFT-based works and RLHF-based works . Hall: Hallucination Rate Resp: response-level hallucination rate, Ment: mention-level hallucination rate. The best result is in bold, and the second-best result is underlined.
Model LLM #Param Object POPE MMHal
HalBench Bench
Resp $\downarrow$ Ment $\downarrow$ F1 $\uparrow$ Score $\uparrow$ Hall $\downarrow$
Teacher MLLM Qwen-1.5-7B 7B 50.1 24.8 84.9 2.60 20.7
Teacher MLLM Qwen-2-7B 7B 29.7 23.4 85.7 2.88 14.5
Qwen-VL-Chat Qwen-7B 9.6B 40.4 20.7 74.9 2.76 38.5
LLaVA-1.5-7B Vicuna-7B 7B 53.6 25.2 86.1 2.36 51.0
VCD Vicuna-1.5-7B 7B 48.8 24.3 84.5 2.12 54.2
OPERA Vicuna-1.5-7B 7B 45.1 22.3 85.4 2.15 54.2
HA-DPO Vicuna-1.5-7B 7B 39.9 19.9 86.8 1.98 60.4
POVID Vicuna-1.5-7B 7B 48.1 24.4 86.3 2.08 56.2
LLaVA-RLHF Vicuna-1.5-13B 13B 38.1 18.9 82.7 2.02 62.5
LURE Vicuna-1.5-7B 7B 27.7 17.3 - 1.64 60.4
RLHF-V Vicuna-13B 13B 12.2 7.5 86.2 2.45 51.0
RLAIF-V Vicuna-1.5-7B 7B 8.5 4.3 - 3.06 29.2
MiniCPM-V MiniCPM-2.4B 2.8B 21.6 11.5 79.5 3.70 24.9
MiniCPM-V-2 MiniCPM-2.4B 2.8B 14.5 7.8 86.3 4.09 18.2
Mini-Gemini-2B Gemma-2B 2B 29.7 21.1 85.6 2.83 18.8
Bunny-2B Qwen-1.5-1.8 2.2B 50.2 23.4 85.8 2.72 19.3
LLaVA-MoD-2B Qwen-1.5-1.8B 2.2B 11.4 7.2 87.0 2.76 17.2
LLaVA-MoD-2B Qwen-2-1.5B 1.9B 11.2 5.9 87.2 2.91 13.8
4.2.2 Hallucination-Oriented Benchmarks
As indicated in Table 3, LLaVA-MoD shows remarkable performance in mitigating hallucination, even beating its teacher model. It can be attributed into two aspects: Firstly, by assigning a higher probability for the positive response, preference distillation encourages the student model to focus on providing correct and relevant information. Secondly, by assigning a lower probability for the negative response, preference distillation discourages incorrect or unsubstantiated information. Using the teacher model as a reference to adjust response probabilities, this optimization enables the student model to handle hallucination problems more accurately and reliably, thereby outperforming the teacher model.
Moreover, LLaVA-MoD surpasses similar small-sized models (Yao et al., 2024; Li et al., 2024c; He et al., 2024) and even recent RLHF-based models (Sun et al., 2023; Zhou et al., 2024b; Huang et al., 2024; Leng et al., 2024; Zhou et al., 2023; Zhao et al., 2023b). For instance, on the Object HalBench, it outperforms RLHF-V (Yu et al., 2024a) by 8.2% in response-level hallucination rate and by 21.3% in mention-level hallucination rate. This demonstrates that taking the teacher model as a reference for preference distillation is a challenging task that can encourage the student model to effectively learn to mitigate hallucination.
4.2.3 Efficiency Comparison
We compare the efficiency of LLaVA-MoD with recent SOTA MLLMs in terms of training data and trainable parameters. As indicated in Table 2, The 2B-sized LLaVA-MoD requires only 5M samples and activates merely a subset of its total 2.2B parameters during training and inference. It demonstrates that LLaVA-MoD achieves 8.8% higher performance than Qwen-VL-Chat, using only 23% trainable parameters and 0.3% of training data samples, respectively. Moreover, when compared to the small-sized model MiniCPM-V with 2.8B trainable parameters, LLaVA-MoD also achieves superior performance with just 1.6% of the data samples and 2B parameters, underscoring its efficiency. In conclusion, LLaVA-MoD enables more efficient utilization of data, parameters, and computational resources. This not only results in faster training and inference but also provides a practical solution for deploying high-performance models in resource-constrained environments.
4.3 Ablation Study
In this section, we begin by analyzing the impact of preference distillation. Next, we conduct ablation studies on a model trained solely with mimic distillation to investigate the effectiveness of our proposed architecture and training strategy on comprehension-oriented benchmarks. We utilize CLIP-ViT-L/14 as the default vision encoder. The default large language models (LLMs) are Qwen-1.5-1.8B and Qwen-1.5-7B for the student and teacher models.
4.3.1 Impact of Preference Distillation
We extensively evaluated preference distillation, examining its impact on both the comprehension ability and hallucination mitigation capabilities of LLaVA-MoD. As detailed in Table 4 and Table 5, preference distillation does not yield consistent improvements in comprehension across all benchmarks, its impact on hallucination mitigation was strikingly consistent. Across all evaluated benchmarks, preference distillation notably reduced the occurrence of hallucinations, highlighting its efficacy in enhancing the reliability and factual accuracy of LLaVA-MoD. This observation aligns with the field of language modeling (Achiam et al., 2023), have consistently indicated that RLHF techniques tend to prioritize the reduction of hallucinations, often at the expense of potential gains in other areas, such as raw comprehension scores.
Table 4: Performance of preference distillation on comprehension-oriented benchmarks.
[…]
Mimic 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
Mimic + Preference 58.7 39.2 68.0 58.5 66.7 66.3 61.9 59.9
Table 5: Performance of preference distillation on hallucination-oriented benchmarks.
Distillation Stage Object POPE MMHal
HalBench Bench
Resp $\downarrow$ Ment $\downarrow$ F1 $\uparrow$ Score $\uparrow$ Hall $\downarrow$
Mimic 39.1 22.6 86.7 2.75 17.8
Mimic + Preference 11.4 7.2 87.0 2.76 17.2
4.3.2 Impact of Training Strategy
Knowledge Distillation Improves Training Efficiency in MoE.
This study explores the impact of knowledge distillation (KD) and examines how it facilitates MoE model training compared to supervised fine-tuning (SFT). The experiments are conducted using the same sparse configuration of E4T2, where 4 experts are initialized and the top-2 experts are activated. The same datasets and training hyperparameters are employed for both KD and SFT.
Training MoE models is inherently challenging, as it requires effectively leveraging huge data to optimize experts’ routing. Our approach demonstrates that KD, compared to traditional SFT, excels in addressing this challenge, achieving superior performance across multiple benchmarks. As demonstrated in Table 6, KD significantly outperforms SFT on all the benchmarks. Specifically, KD achieves an average improvement of 8.1% over SFT, with particularly pronounced gains in complex multi-task scenarios. For instance, on the MMB and MME benchmarks, KD surpasses SFT by 8.2% and 10.0%, respectively. These results highlight KD’s ability to effectively leverage multiple experts and extract comprehensive knowledge from large-scale MLLMs. By distilling essential knowledge, KD not only maximizes data utilization but also achieves better generalization ability of MoE compared to SFT.
Table 6: Comparison between the KD and SFT. The architecture of the student model is configured as E4T2.
[…]
SFT 58.6 31.2 66.2 55.6 63.2 59.2 56.1 55.7
KD 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
Focusing on Response Improves Distillation Generalization.
This study explores the impact of the distillation target. The default approach distills response, we experiment with distilling both response and instruction. Both settings utilize the same datasets and training hyperparameters. Table 7 shows that distilling only the response achieves superior performance, surpassing jointly distilling response and instruction by 3.1% in terms of average performance. This suggests that focusing on response is more effective for knowledge distillation in auto-regressive modeling. A possible explanation lies in that mimicking instructions could lead to overfitting the specific instruction patterns of training data, potentially reducing the model’s adaptability to unseen instructions. Mimicking the teacher model’s response proves to be effective in improving the student’s auto-regressive modeling generalization capabilities.
Table 7: Comparison between different distillation tokens. The response indicates distilling solely on the generated response tokens, while the response+instruction incorporates additional instruction tokens in the distillation process.
[…]
Response + Instruction 58.5 35.0 66.8 58.3 68.5 61.8 59.0 58.4
Response 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
4.3.3 Impact of Model Architecture
Sparse MoE Enables Effective Knowledge Distillation.
This study explores the effect of sparse architecture in distillation by comparing its performance to dense architecture. We adopt the E4T1 configuration as the sparse architecture to ensure the same activated parameters as the dense architecture. The same datasets and training hyperparameters are used for sparse and dense architectures.
[…]
Table 8: Comparison between sparse and dense architecture within the distillation. The sparse architecture is configured as E4T1.
[…]
Dense 57.6 32.7 67.3 56.8 65.2 61.8 58.2 57.1
Sparse 58.7 36.9 67.9 58.2 66.1 64.5 61.7 59.2
Teacher Capacity Matters in Knowledge Distillation.
This study explores the impact of the teacher in the distillation process. Specifically, we utilize CLIP-ViT-L/14 as the vision encoder. For the LLM, Qwen-1.5-7B is the strong teacher while Qwen-1.5-4B is the weaker teacher. All experiments are conducted under the same datasets and training hyperparameters to ensure a fair comparison.
As shown in Table 9, a suitable teacher enhances performance across various benchmarks. For instance, the Qwen-1.5-1.8B student model achieves an average performance of 62.3% with the weaker Qwen-1.5-4B teacher. This performance jumps to 63.6% when utilizing the stronger Qwen-1.5-7B teacher. Interestingly, the smaller Qwen-1.5-0.5B student model shows marginal improvement with either the Qwen-1.5-4B or Qwen-1.5-7B teacher. This suggests that an excessively large capacity gap between the teacher and student models can hinder effective knowledge transfer. Utilizing a ”middle teacher” with an intermediate capacity could bridge this gap, facilitating smoother knowledge transfer and boosting the student model’s learning efficiency.
Table 9: Comparison between the strong and weak teachers within the distillation. We set the strong teacher as the LLM is Qwen-1.5-7B and the weak teacher as the LLM is Qwen-1.5-4B. The model configuration is E4T2.
[…]
Qwen-1.5-0.5B Qwen-1.5-4B 56.0 25.3 64.7 53.8 63.3 62.2 50.8 53.7
Qwen-1.5-7B 56.1 29.8 63.8 54.5 64.2 58.5 49.7 53.8
Qwen-1.5-1.8B Qwen-1.5-4B 58.7 34.6 67.9 57.7 67.6 64.9 60.7 58.9
Qwen-1.5-7B 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
5 Conclusion
In this paper, we introduce LLaVA-MoD, a novel framework for efficient training of small-scale multimodal language models via knowledge distillation from large-scale ones. It addresses two key challenges in MLLM distillation: enhancing s-MLLM architecture with MoE design for efficiency and expressiveness balance and implementing a progressive knowledge transfer strategy. Extensive experiments show that LLaVA-MoD outperforms existing models with low activation parameters and computational costs. Notably, it outperforms Qwen-VL-Chat-7B by 8.8% with only 2 billion activation parameters, 0.3% training data, and 23% trainable parameters, highlighting its effectiveness in distilling knowledge and driving more efficient multimodal language model development.
6 Limitations
LLaVA-MoD requires that both the teacher and student models belong to the same family of large language models (LLMs) to ensure consistency in the vocabulary space. Future research could explore distillation techniques involving heterogeneous model families. Additionally, the need to load both student and teacher models results in significant memory usage. To facilitate more efficient distillation, a feasible solution would be to pre-extract the outputs of the teacher model, allowing only the student model to be loaded during training.
Appendix A Appendix
A.1 Dataset
We report the detailed training dataset settings of LLaVA-MoD in Table 10.
Table 10: Training dataset. #Sample means the training samples.
Stage Task Dataset
Initialization Captioning LLaVA-1.5-Pretrain (Liu et al., 2023b)
Dense-to-Dense Distillation Captioning ALLaVA-Caption-4V (Chen et al., 2024), ShareGPT4V-PT (Chen et al., 2023a)
Conversation MIMIC-IT (Li et al., 2023a), LVIS (Wang et al., 2023a), LRV (Liu et al., 2023a), SViT (Zhao et al., 2023a)
Dense-to-Sparse Distillation Captioning ShareGPT4V-100K (Chen et al., 2023a), TextCaps (Sidorov et al., 2020)
Conversation LLaVA-1.5-Instruct (Liu et al., 2023b)
General QA GQA (Hudson & Manning, 2019), VQAv2 (Goyal et al., 2017), OKVQA (Marino et al., 2019)
Grounding Visual Genome (Krishna et al., 2017), RefCOCO (Yu et al., 2016; Mao et al., 2016)
Science AI2D (Kembhavi et al., 2016), ScienceQA (Lu et al., 2022a)
Chart & Doc DVQA (Kafle et al., 2018), ChartQA (Masry et al., 2022), DocQA (Clark & Gardner, 2018)
OCR OCRVQA (Mishra et al., 2019), SynthDoG-EN (Kim et al., 2022)
Knowledge A-OKVQA (Schwenk et al., 2022), GeoQA+ (Cao & Xiao, 2022)
Preference Distillation Preference RLAIF-V (Yu et al., 2024b)
A.2 Hyperparameters
We report the detailed training hyperparameters of LLaVA-MoD in Table 11.
Table 11: Training hyperparameters
Configuration Stage-1 Stage-2 Stage-3
LLM
VL Adaptor
ViT
LLM init. Qwen-1.5-1.8B Qwen-1.5-1.8B 2nd-stage
VL Adaptor init. MLP 1st-stage 2nd-stage
ViT init. CLIP-Large@336
Image resolution 336 $\times$ 336
ViT sequence length 576
LLM sequence length 2048
Optimizer AdamW
Optimizer hyperparameter $\beta_{1}=0.9,\beta_{2}=0.98,eps=1e^{-6}$
Learning rate 1e-4 2e-5 2e-5
Learning rate schedule Cosine decay
Weight decay 0.0
Training epoch 1
Warm-up ratio 0.03
Global batch size 256 128 128
Numerical precision Bfloat16
Model parallelism Zero2 Zero2 offload Zero2 offload

### v3

@@
 Abstract
-We introduce LLaVA-MoD, a novel framework designed to enable the efficient training of small-scale Multimodal Language Models (s-MLLM) by distilling knowledge from large-scale MLLM (l-MLLM).
+We introduce LLaVA-MoD, a novel framework designed to enable the efficient training of small-scale Multimodal Language Models (s-MLLM) distilling knowledge from large-scale MLLM (l-MLLM).
 Our approach tackles two fundamental challenges in MLLM distillation.
 First, we optimize the network structure of s-MLLM by integrating a sparse Mixture of Experts (MoE) architecture into the language model, striking a balance between computational efficiency and model expressiveness.
-Second, we propose a progressive knowledge transfer strategy to ensure comprehensive knowledge migration.
-This strategy begins with mimic distillation, where we minimize the Kullback-Leibler (KL) divergence between output distributions to enable the student model to emulate the teacher network’s understanding.
-Following this, we introduce preference distillation via Direct Preference Optimization (DPO), where the key lies in treating l-MLLM as the reference model.
-During this phase, the s-MLLM’s ability to discriminate between superior and inferior examples is significantly enhanced beyond l-MLLM, leading to a better student that surpasses its teacher, particularly in hallucination benchmarks.
-Extensive experiments demonstrate that LLaVA-MoD outperforms existing models across various multimodal benchmarks while maintaining a minimal number of activated parameters and low computational costs.
-Remarkably, LLaVA-MoD, with only 2B activated parameters, surpasses Qwen-VL-Chat-7B by an average of 8.8% across benchmarks, using merely $0.3\%$ of the training data and 23% trainable parameters.
-These results underscore LLaVA-MoD’s ability to effectively distill comprehensive knowledge from its teacher model, paving the way for the development of more efficient MLLMs.
-The code will be available on https://github.com/shufangxun/LLaVA-MoD
+Second, we propose a progressive knowledge transfer strategy for comprehensive knowledge transfer.
+This strategy begins with mimic distillation, where we minimize the Kullback-Leibler (KL) divergence between output distributions to enable s-MLLM to emulate l-MLLM’s understanding.
+Following this, we introduce preference distillation via Preference Optimization (PO), where the key lies in treating l-MLLM as the reference model.
+During this phase, the s-MLLM’s ability to discriminate between superior and inferior examples is significantly enhanced beyond l-MLLM, leading to a better s-MLLM that surpasses l-MLLM, particularly in hallucination benchmarks.
+Extensive experiments demonstrate that LLaVA-MoD surpasses existing works across various benchmarks while maintaining minimal activated parameters and low computational costs.
+Remarkably, LLaVA-MoD-2B surpasses Qwen-VL-Chat-7B with an average gain of 8.8%, using merely $0.3\%$ of the training data and 23% trainable parameters.
+The results underscore LLaVA-MoD’s ability to effectively distill comprehensive knowledge from its teacher model, paving the way for developing efficient MLLMs.
+The code is available at https://github.com/shufangxun/LLaVA-MoD.
+1 Introduction
+Multimodal Large Language Models (MLLM) (Bai et al., 2023b; Liu et al., 2024; Lin et al., 2024b; Chen et al., 2023b; Shu et al., 2023; Lu et al., 2024) have shown promising results in multimodal tasks by integrating visual encoders with Large Language Models (LLM) (Achiam et al., 2023; Bai et al., 2023a; Team et al., 2023; Touvron et al., 2023).
+However, the large model size and extensive training data present significant computational challenges.
+For example, the largest version of LLaVA-NeXT employs Qwen-1.5-110B and trains with 128 H800 GPUs for 18 hours.
+Additionally, the substantial number of parameters requires advanced hardware and results in slow inference speed, complicating real-world deployment, especially on mobile devices.
+Therefore, exploring a small-scale MLLM (s-MLLM) that balances performance and efficiency is a crucial topic.
 Figure 1: Comparisons of training cost and performance.
-LLaVA-MoD achieves competitive performance with advanced MLLMs using significantly lower training costs while outperforming current small-scale MLLMs by a large margin.
-1 Introduction
-Leveraging the advanced instruction-following and reasoning capabilities of Large Language Models (LLMs) (Achiam et al., 2023; Bai et al., 2023a; Jiang et al., 2024; Team et al., 2023; Touvron et al., 2023a; b; Dubey et al., 2024), Multimodal Large Language Models (MLLMs) (Bai et al., 2023b; Liu et al., 2024; Lin et al., 2024b; Li et al., 2023b; Chen et al., 2023b; Shu et al., 2023; Lu et al., 2024), typically integrate a visual encoder (Caron et al., 2021; Radford et al., 2021; Liu et al., 2022) alongside an LLM to achieve promising results across various vision and cross-modal tasks, such as image captioning and visual question-answering.
-MLLMs are characterized by their large model size and extensive training data, which contribute to their superior performance but also necessitate significant computation resources.
-For instance, the largest version of LLaVA-NeXT (Li et al., 2024a) employs Qwen-1.5-110B (Yang et al., 2024) as the language model and requires 18 hours of training with 128 H800 GPUs.
-Furthermore, MLLMs face deployment challenges due to the high memory and computational demands.
-For instance, models with substantial parameters often require advanced hardware and exhibit slow inference speeds, which significantly hinder their real-world applications such as on mobile devices.
-Thus, developing MLLMs that balance performance and efficiency has become a crucial research focus.
-Previous works on small-scale Multimodal Large Language Models (s-MLLM) (Zhou et al., 2024a; Yuan et al., 2023; Shao et al., 2024; He et al., 2024; Chu et al., 2023; 2024; Yao et al., 2024) have mainly concentrated on devising data collection (Schuhmann et al., 2021; 2022; Zhu et al., 2024; Awadalla et al., 2024; Gadre et al., 2024; Chen et al., 2023a) and filtering protocols (Xu et al., 2023; Fan et al., 2024; Zhang et al., 2023; Li et al., 2024b; Nguyen et al., 2024) to ensure high-quality training data.
-While such high-quality data aid s-MLLM in narrowing the performance gap with their larger counterparts, their effectiveness is still limited by model capacity and the availability of high-quality data.
-With the advent of open-source MLLMs, an intuitive strategy inspired by conventional deep learning tasks involves utilizing large-scale MLLMs (l-MLLM) as the teacher to improve the training of s-MLLM via Knowledge Distillation (KD) (Hinton et al., 2015).
-KD facilitates stable and efficient training of smaller models by aligning the output distributions of the student model with those of the teacher model.
-However, the application of KD as a model reduction technique has yet to be extensively explored in MLLMs.
-To develop an effective framework for MLLM distillation, we consider two primary challenges.
-The first challenge lies in designing a lightweight architecture for student MLLMs that retains strong learning and expressive capabilities, enabling it to absorb the complex knowledge embedded within teacher MLLMs effectively.
-The second challenge is to efficiently and comprehensively transfer this knowledge from the teacher MLLMs to the student MLLMs.
-To address these two challenges, we present LLaVA-MoD [^1] MoD denotes Mixture-of-Expert Knowledge Distillation for MLLM, an effective framework for training s-MLLM by mimicking the behavior of the l-MLLM through Mixture-of-Expert (MoE) Knowledge Distillation.
-For the first challenge—designing a lightweight s-MLLM structure—a straightforward approach might involve reducing the scale of the base language model in l-MLLM to create a smaller network.
-However, this direct reduction in network parameters can significantly compromise the model’s expressive capability, making it less effective in modeling complex multimodal understanding tasks.
-Drawing inspiration from the recent success of sparse MoE (Lin et al., 2024a; Dai et al., 2024; Jiang et al., 2024; Shen et al., 2023) in sequential modeling, we incorporate a sparse MoE structure into the dense s-MLLM.
-This approach seeks to balance scale reduction while preserving the model’s ability to capture and represent intricate multimodal information during distillation.
-Specifically, we enhance the s-MLLM by equipping it with several feedforward networks (FFNs) and linear gates within the LLM.
-Each FFN acts as an expert, capturing fine-grained knowledge from the l-MLLM, while the gates select the top-$k$ experts to facilitate optimal knowledge transfer pathways.
-To address the second challenge, we propose a progressive distillation strategy for effective knowledge transfer.
-The process begins by aligning the vision encoder with the LLM using a learnable adapter, thereby initializing a dense student MLLM.
-Following this, we employ two consecutive distillation stages, where the student model evolves from mimicking and approximating the teacher MLLM to ultimately surpassing it: Mimic Distillation.
-This stage is divided into two steps, i.e., dense-to-dense and dense-to-sparse distillation.
-In the dense-to-dense distillation stage, we employ standard knowledge distillation loss to align the output logits distribution between the initialized dense s-MLLM and the teacher model, using general captioning and conversation datasets.
-Next, we transition to the dense-to-sparse distillation where we transform the dense student MLLM into a sparse one by integrating MoE, and then distill knowledge from the teacher MLLM into this sparse student model using a broad range of tasks and datasets.
+LLaVA-MoD achieves comparable performance with advanced MLLMs using significantly lower training costs while outperforming current small-scale MLLMs by a large margin.
+Previous works on s-MLLM (Zhou et al., 2024a; Yuan et al., 2023; Shao et al., 2024; He et al., 2024; Chu et al., 2023; Yao et al., 2024) have focused on high-quality data collection and filtering protocols.
+While effective, they are inherently limited by the model capacity.
+Knowledge Distillation (KD) from large-scale MLLM (l-MLLM) offers a promising yet unexplored strategy for enhancing s-MLLM performance.
+By aligning small models’ output distributions with large models, KD allows s-MLLM to leverage the rich knowledge embedded in l-MLLM.
+We explore a distillation framework for MLLM and consider two primary challenges.
+The first challenge is to design a lightweight architecture for s-MLLM while still maintaining strong learning and expressive capabilities to effectively absorb complex knowledge from l-MLLM.
+The second challenge is to ensure an effective transfer of complex, multi-task knowledge from l-MLLM to s-MLLM.
+We present LLaVA-MoD [^1] MoD denotes Mixture-of-Expert Knowledge Distillation for MLLMs to address the challenges through Mixture-of-Expert (MoE) and Knowledge Distillation.
+To address the first challenge, one straightforward approach is to create a small model by simply scaling down l-MLLM.
+However, this direct reduction in size can significantly diminish the model’s expressive capabilities, resulting in a decline in performance for handling complex multimodal tasks.
+Drawing inspiration from the recent success of MoE (Dai et al., 2024; Jiang et al., 2024) in language modeling, We design an MoE structure into s-MLLM to balance the scale reduction while maintaining the ability to capture complex multimodal knowledge through sparsely activated experts during distillation.
+Specifically, we equip s-MLLM with multiple feedforward networks (FFNs) and a linear gate within the LLM block.
+Each FFN serves as an expert to capture fine-grained knowledge from l-MLLM, while the gate dynamically selects top-$k$ experts to identify the optimal knowledge transfer pathway and maintain the training and inference efficiency.
+To address the second challenge, we propose a progressive distillation strategy.
+The process begins by aligning the vision encoder with the LLM using a learnable adapter, thereby initializing a dense s-MLLM.
+Subsequently, we employ two consecutive distillation stages, where s-MLLM evolves from mimicking and approximating l-MLLM to ultimately surpassing it: Mimic Distillation.
+This stage is divided into two steps, i.e., dense-to-dense (D2D) and dense-to-sparse (D2S).
+The rationale for this two-step process is to facilitate the transfer of complex knowledge, where the first step targets general knowledge to build a solid foundation, while the second step targets specialized knowledge to handle multi-task tasks.
+In D2D, we employ standard KD loss to align the output logits distribution between the initialized dense s-MLLM and l-MLLM, using general captioning and conversation datasets.
+Next, in D2S, s-MLLM is first transformed from dense to sparse.
+We then employ the standard KD loss and LM loss to distill the MoE block, using the complex multi-task datasets.
 Preference Distillation.
-In this stage, the teacher model provides knowledge regarding what constitutes “good” and “bad” samples, establishing a foundational reference for the student model.
-The student MLLM leverages this knowledge to adjust its probability distribution, ensuring that good samples have a higher probability than those from the teacher model, while bad samples are assigned a lower probability.
-This process enhances the student model’s ability to mitigate hallucinations by improving its judgment capabilities beyond those of the teacher model.
-As illustrated in Figure 1, LLaVA-MoD exhibits impressive performance across various multimodal benchmarks while maintaining minimal activated parameters and low computational resources.
-For example, LLaVA-MoD-2B exceeds Qwen-VL-Chat-7B by an average of $8.8\%$ on these benchmarks, utilizing just $0.3\%$ of the training data and $23\%$ trainable parameters.
+In this stage, l-MLLM provides knowledge regarding what constitutes “good” and “bad” samples, establishing a foundational reference for the student model.
+The s-MLLM leverages this knowledge to adjust its probability distribution, ensuring that good samples have a higher probability than those from l-MLLM, while bad samples are assigned a lower probability.
+This process enhances the s-MLLM’s ability to mitigate hallucinations by improving its judgment capabilities beyond those of l-MLLM.
+LLaVA-MoD exhibits impressive performance across various multimodal benchmarks while maintaining minimal activated parameters and low computational resources.
+As illustrated in Figure 1, LLaVA-MoD-2B exceeds Qwen-VL-Chat-7B by an average of $8.8\%$ on these benchmarks, utilizing just $0.3\%$ of the training data and $23\%$ trainable parameters.
 Furthermore, it matches the performance of RLHF-based methods with 7B and 13B parameters on several hallucination benchmarks.
 Specifically, LLaVA-MoD-2B surpasses RLHF-V (Yu et al., 2024a) by $8.2\%$ in response-level hallucination rate and by $21.3\%$ in mention-level hallucination rate on the Object HalBench.
-The impressive results demonstrate the effectiveness of our proposed MoD framework in transferring knowledge from l-MLLM to s-MLLM.
+The impressive results demonstrate the effectiveness of our distillation framework in transferring knowledge from l-MLLM to s-MLLM.
 2 Related Work
 Multimodal Large Language Models.
-The emergence of large language models (LLMs) has greatly advanced the field of natural language processing.
-Connecting visual information to LLM to enhance their understanding of multimodal input is crucial for promoting the unified comprehension of vision and language.
-CLIP (Radford et al., 2021) was the first to align visual and textual information into a unified embedding space through contrastive learning objectives.
-BLIP-2 (Li et al., 2023b) utilizes various pre-training tasks to train separate visual and text streams, and adds additional intermediary structures to adapt visual features to LLM.
-Flamingo (Alayrac et al., 2022) incorporates additional cross-attention modules into LLM to handle arbitrarily interleaved sequences of images and texts.
-Recently, models like LLaVA (Liu et al., 2024) and MiniGPT-4 (Zhu et al., 2023) use linear projection layers to map image features to the textual space and enhance the models’ instruction-following ability through visual instruction tuning.
-Moreover, some works focus on a stronger vision encoder.
-Intern-VL (Chen et al., 2023b) leverages a larger-scale vision encoder to bridge the gap with a large-scale language model, and Mini-Gemini (Li et al., 2024c) introduces additional vision encoders for high-resolution refinement to enhance visual tokens.
-Qwen-VL (Bai et al., 2023b), VisionLLM (Wang et al., 2023c), and others aim to imbue models with stronger fine-grained visual understanding capabilities, such as grounding and region understanding.
-Unlike these approaches, our method does not emphasize enlarging the model to boost multimodal capabilities.
-Instead, we adopt a distillation approach to compress the capabilities of larger models into a smaller, sparse MoE architecture, enhancing computational and storage efficiency.
+LLMs have greatly advanced the field of natural language processing.
+Connecting visual information to LLM is crucial for promoting the unified comprehension of vision and language.
+BLIP-2 (Li et al., 2023b) adds additional intermediary structures to adapt visual features to LLM.
+Flamingo (Alayrac et al., 2022) incorporates additional cross-attention modules into LLM to handle arbitrarily interleaved multimodal sequences.
+Recently, models like LLaVA (Liu et al., 2024) and MiniGPT-4 (Zhu et al., 2023) have tried to enhance the models’ instruction-following ability through visual instruction tuning.
+In addition, some works focus on a stronger vision encoder (Chen et al., 2023b; Li et al., 2024) or more powerful fine-grained visual understanding capabilities (Bai et al., 2023b; Wang et al., 2024).
+Unlike these approaches, our method does not need to significantly enlarge the model.
+We aim to combine distillation and MoE to improve computational and storage efficiency while maintaining advanced performance.
 Knowledge Distillation.
-Large language models possess powerful capabilities, but their enormous size and high inference costs limit their application in low-resource environments.
-Knowledge distillation (Hinton et al., 2015) uses a large model as teacher to transfer its advanced knowledge to a smaller student model, which plays a critical role in compressing model size and enables smaller models to self-improve.
-Some works adopt additional designs to make the distillation process more suitable for LLMs.
-MiniLLM (Gu et al., 2023) minimizes the reverse KLD (Kullback–Leibler divergence) to prevent students from overestimating low-probability regions in the teacher distribution, while GKD (Agarwal et al., 2023) introduces generalized knowledge distillation and promotes the integration of distillation with RLHF (Reinforcement Learning from Human Feedback).
-Additionally, Huang et al.
-(Huang et al., 2022) introduce context distillation, and some works (Mukherjee et al., 2023; Li et al., 2022; Ho et al., 2023) adopt CoT (Chain-of-Thought) distillation to enhance specific skills of small models.
-Our approach innovatively distills the knowledge of MLLM into a smaller sparse MoE (Mixture-of-Experts) architecture, significantly enhancing the multimodal processing capabilities of small models at a minimal cost.
-Mixture-of-Experts The mixture of experts (MoE) architecture is initially introduced by Jacobs et al.
-(1991) to enhance model performance across various samples through the utilization of independent experts.
-Recently, with the rise of transformer architectures, numerous studies have extended transformer layers to incorporate MoE.
-In transformer-based MoE architecture, feed-forward neural network (FFN) layers are often replaced with sparsely activated experts, employing a trainable Top-k gating strategy (Lepikhin et al., 2020; Fedus et al., 2022), thereby effectively increasing model capacity while maintaining a low computational overhead.
-Additionally, to further mitigate training costs, the sparse upcycling method (Komatsuzaki et al., 2022) is adopted, which initializes expert parameters using those from a well-trained dense model.
-Currently, MoE architectures have found widespread application not only in pre-trained language models (Jiang et al., 2024; Dai et al., 2024) but also in vision models and vision-language models (Lin et al., 2024a; Shen et al., 2023).
-Our approach integrates MoE with knowledge distillation techniques to provide stronger signals for sparse training, which remarkably decrease the training cost associated with sparse models.
+The enormous sizes and high inference costs of LLMs limit their application in low-resource environments.
+Knowledge distillation (Hinton et al., 2015) uses a large model as the teacher to transfer its advanced knowledge to a smaller student model, which plays a critical role in compressing model size and enables smaller models to self-improve.
+MiniLLM (Gu et al., 2023) minimizes the reverse KL divergence to prevent students from overestimating low-probability regions in the teacher distribution, while GKD (Agarwal et al., 2023) introduces generalized knowledge distillation and promotes the integration of distillation with RLHF.
+Additionally, some works adopt context distillation (Huang et al., 2022) and CoT (Chain-of-Thought) distillation (Mukherjee et al., 2023; Li et al., 2022; Ho et al., 2022) to enhance specific skills of small models.
+Our approach innovatively distills the knowledge of MLLM into a smaller sparse MoE architecture, significantly enhancing the multimodal processing capabilities of small models at a minimal cost.
+Mixture-of-Experts The mixture of experts (MoE) architecture, introduced by (Jacobs et al., 1991), enhances performance by leveraging independent experts to handle diverse samples.
+In transformer-based architecture, the feed-forward neural network (FFN) layers serve as the experts, sparsely activated through a gating strategy (Lepikhin et al., 2020; Fedus et al., 2022).
+This design effectively enhances model capacity while keeping computational overhead low.
+Additionally, the sparse upcycling method (Komatsuzaki et al., 2022) which initializes experts using those from a well-trained dense model is proposed to further mitigate training costs.
+MoE has shown great potential not only in language models (Jiang et al., 2024; Dai et al., 2024) but also in vision models (Riquelme et al., 2021) and vision-language models (Lin et al., 2024a; Shen et al., 2023).
+Our approach integrates MoE with knowledge distillation techniques to provide stronger signals for sparse training, which remarkably decreases the training cost associated with sparse models.
 3 Method
@@
 Preference Distillation: Increasing the student’s positive response probabilities $p^{+}_{S}$ to surpass those of the teacher $p^{+}_{T}$, while decreasing the student’s negative response probabilities $p^{-}_{S}$ to fall below those of the teacher $p^{-}_{T}$, via the Preference-Optimization (PO) loss.
-We introduce LLaVA-MoD, a novel framework for building efficient MLLM using mixture-of-experts (MoE) knowledge distillation.
+We introduce LLaVA-MoD, a novel framework for building efficient s-MLLM using mixture-of-experts (MoE) and knowledge distillation.
 Our framework consists of two main components: (a).
-Architecture Design of s-MLLM: As shown in Figure 3, we design the s-MLLM with a sparse MoE framework, enhancing the ability to acquire specialized expert knowledge.
+Architecture Design of s-MLLM: As shown in Fig.
+3, we design a sparse s-MLLM with MoE, enhancing the ability to acquire specialized expert knowledge while maintaining training and inference efficiency.
 (b).
-Distillation Mechanism: We design a progressive distillation mechanism as shown in Figure 2 to transfer knowledge from the l-MLLM to the sparse s-MLLM.
+Distillation Mechanism: We design a progressive distillation mechanism as shown in Fig.
+2 to transfer knowledge from l-MLLM to sparse s-MLLM.
 This process involves two stages: mimic distillation followed by preference distillation.
-We will introduce LLaVA-MoD in detail from these two components.
 3.1 Architecture Design of Sparse s-MLLM
@@
 s-MLLM Definition.
-As illustrated in Figure 3, the basic architecture of s-MLLM consists of three primary components: a vision encoder, a large language model (LLM), and a vision-language (VL) adaptor.
+As illustrated in Fig.
+3, the basic architecture of s-MLLM consists of three primary components: a vision encoder, a large language model (LLM), and a vision-language (VL) adaptor.
 Given a multimodal instruction conversation $(x,y)$, we define our s-MLLM to process response $y$ as follows:
@@
 $\text{LLM}_{\phi}$ is the large language model with parameters $\phi$, which produces the response $y$ based on the multimodal tokens of $x=[x_{v},x_{i}]$.
-Figure 3: Dense to Sparse Architecture of LLaVA-MoD.
-The architecture of LLaVA-MoD consists of three components: VL Adaptor, Vision Encoder, and LLM.
-The VL Adaptor and Vision Encoder remain unchanged, while the LLM is upcycled from dense to sparse.
-Mixture-of-Experts.
+Figure 3: Sparsification of s-MLLM. The VL Adaptor and Vision Encoder remain unchanged, while the LLM is upcycled from dense to sparse.
+Sparsify s-MLLM.
 The principle of building our s-MLLM is downsizing the LLM while leaving the vision encoder and vision-language adaptor unmodified.
 To achieve this downsizing goal, we sparsify the dense s-MLLM by incorporating an MoE architecture.
-Specifically, Figure 3 illustrates the process, where we apply the sparse upcycling technique (Komatsuzaki et al., 2022) to replicate $N$ feedforward networks (FFNs) as the expert modules.
+Specifically, Fig.
+3 illustrates the process, where we apply the sparse upcycling technique (Komatsuzaki et al., 2022) to replicate $N$ feedforward networks (FFNs) as the expert modules.
 Additionally, we introduce a linear layer as the router, which dynamically activates the appropriate experts by predicting the probability of expert assignment.
@@
 3.2 Progressive Distillation
-Our progressive distillation consists of two distinct stages, i.e., mimic distillation (Figure 2 (a)) and preference distillation (Figure 2 (b)).
-In the mimic distillation stage, the student MLLM $\pi_{\text{S}}$ imitates the general and specific knowledge from the teacher MLLM $\pi_{\text{T}}$.
-In the preference distillation stage, the student MLLM gains the preference knowledge of the teacher MLLM to further refine its output and reduce hallucinations.
-We use the pre-trained l-MLLM as the teacher MLLM, which is frozen during the whole distillation process.
-Both teacher and student MLLMs are in the same LLM family, ensuring consistent vocabulary space, which is essential for the student MLLM to imitate the teacher MLLM accurately.
-For simplicity, we denote the student as $\pi_{\text{S}}$ and the teacher as $\pi_{\text{T}}$.
+Our progressive distillation consists of two distinct stages, i.e., mimic distillation (Fig.
+2 (a)) and preference distillation (Fig.
+2 (b)).
+In the mimic distillation stage, the s-MLLM $\pi_{\text{S}}$ imitates the general and specific knowledge from the l-MLLM $\pi_{\text{T}}$.
+In the preference distillation stage, $\pi_{\text{S}}$ gains the preference knowledge of $\pi_{\text{T}}$ to further refine its output and reduce hallucinations.
+Both $\pi_{\text{S}}$ and $\pi_{\text{T}}$ are from the same LLM family.
+This ensures a consistent vocabulary space, which is essential for accurate imitation.
 Initialization.
-Before distillation, we first align the vision encoder with the LLM via a learnable adapter, aiming to obtain a well-initialized dense student MLLM.
+Before distillation, we first align the vision encoder with the LLM via a learnable adapter, aiming to obtain a well-initialized dense version of $\pi_{\text{S}}$.
 The $\text{LLM}_{\phi}$ and $\text{ViT}_{\chi}$ are maintained frozen since their pre-trained parameters have already captured rich visual and language knowledge.
@@
 The training objective is to minimize the cross-entropy of the generated tokens.
-The objective function is defined as:
+The objective function is:
 $\mathcal{L}_{\text{Init}}(\pi_{\text{S}})=-\mathbb{E}_{(y_{k}\mid y_{<k},x)\sim\pi_{\text{S}}}\left[\log\pi_{\text{S}}(y_{k}\mid y_{<k},x)\right],$ (5)
@@
 Mimic Distillation.
-We decompose the comprehensive knowledge of the teacher MLLM into general and expertise aspects and undertake dense-to-dense distillation and dense-to-sparse distillation respectively to transfer this knowledge to the student MLLM.
+We decompose the comprehensive knowledge within $\pi_{\text{T}}$ into general and specialized aspects to address the challenges posed by their structural differences, which can complicate simultaneous learning.
+Subsequently, we conduct a general-to-specialized mimic distillation, which consists of two steps: dense-to-dense (D2D) and dense-to-sparse (D2S) distillation, to transfer the knowledge to $\pi_{\text{S}}$.
+This two-step approach balances the transfer of general and specialized knowledge through the progressive distillation, thereby enhancing overall performance.
+As illustrated in Fig.
+3, we utilize the dense structure of $\pi_{\text{S}}$ during D2D to acquire general knowledge and transform it into a sparse structure during D2S to acquire complex specialized knowledge.
+Throughout both stages, $\pi_{\text{T}}$ remains unchanged.
 a).
 Dense-to-Dense Distillation.
-In this stage, we aim to replicate the general knowledge of the teacher MLLM.
-Acquiring general knowledge first is crucial, as it provides a broad foundation and common understanding across various domains, enabling the student MLLM to develop a basic framework applicable in multiple scenarios.
-This foundation supports a more comprehensive and flexible understanding before progressing to specialized tasks.
+In this stage, we aim to replicate the general knowledge of l-MLLM.
+Acquiring general knowledge first is crucial, as it establishes a broad foundation across various scenarios, enabling s-MLLM to develop a basic framework before advancing to multiple specialized tasks.
 To achieve this, we maintain $\text{ViT}_{\chi}$ frozen, and jointly optimize $\text{LLM}_{\phi}$ and $\text{Proj}_{\omega}$, with trainable parameters $\theta=\{\omega,\phi\}$.
 We leverage common image-caption pairs and conversation datasets.
-The training objective is to minimize the Kullback-Leibler divergence (KLD) between the output logits of the dense student MLLM and the teacher MLLM.
-The objective function is defined as:
+The training objective is to minimize the Kullback-Leibler divergence (KLD) between the output logits of s-MLLM and l-MLLM.
+The objective function is:
 $\mathcal{L}_{\text{D2D}}(\pi_{\text{S}};\pi_{\text{T}})=-\mathbb{E}_{(x,y_{k})\sim\pi_{\text{T}}}\left[\log\frac{\pi_{\text{T}}(y_{k}\mid y_{<k},x)}{\pi_{\text{S}}(y_{k}\mid y_{<k},x)}\right],$ (6)
-where $V$ denotes the vocabulary, while $\pi_{\text{S}}(y_{k}\mid y_{<k},x)$ and $\pi_{\text{T}}(y_{k}\mid y_{<k},x)$ denote the probability of the predicted tokens for student and teacher models, respectively.
+where $V$ denotes the vocabulary, while $\pi_{\text{S}}(y_{k}\mid y_{<k},x)$ and $\pi_{\text{T}}(y_{k}\mid y_{<k},x)$ denote the probability of the predicted tokens for s-MLLM and l-MLLM.
 b).
 Dense-to-Sparse Distillation.
-In this stage, our focus shifts to transferring the specialized expertise of the teacher MLLM, allowing the student MLLM to obtain advanced capabilities and achieve superior performance in diverse tasks.
-However, directly learning this knowledge in its dense form could lead to inefficiencies and difficulty.
-Therefore, we sparsify the dense student MLLM by introducing multiple experts.
-As detailed in Section 3.1, we replicate $N$ feedforward networks (FFNs) within $\text{LLM}_{\phi}$ and incorporate an MLP layer as a router, which jointly constitutes the experts with parameters $\phi_{e}$.
-This sparse architecture allows the student MLLM to selectively activate the most relevant experts for different tasks and inputs, offering significant advantages in emulating the teacher’s specialized knowledge.
+In this stage, our focus shifts to transfer the specialized knowledge of l-MLLM into s-MLLM to obtain advanced capabilities and achieve superior performance in complex tasks.
+However, directly learning this knowledge in the dense form of s-MLLM could be insufficient due to the diverse knowledge structure of different tasks.
+Therefore, we sparsify the dense s-MLLM by introducing multiple experts.
+As detailed in Section 3.1, we replicate $N$ feedforward networks (FFNs) within $\text{LLM}_{\phi}$ and add an MLP layer as a router, forming the experts with parameters $\phi_{e}$.
+This sparse architecture enables s-MLLM to activate the most relevant experts based on different inputs selectively, offering significant advantages in emulating the specialized knowledge of l-MLLM.
 For training, we leverage multi-task data, updating only the experts and the adaptor.
@@
 The trainable parameters are $\theta=\{\omega,\phi_{e}\}$.
-Similar to the previous stage, we also adopt the KLD as the training objective.
-Additionally, we include the standard next-token prediction objective, i.e., minimizing the student-generated tokens’ cross-entropy loss, to inject supervision from ground truth data which can reduce the existing bias in the teacher model.
-The final objective function is defined as:
+Similar to the previous stage, we adopt the KLD as the training objective.
+Additionally, we include the standard next-token prediction objective, i.e., minimizing the cross-entropy loss of generated tokens by s-MLLM, to inject supervision from ground truth data which can reduce the existing bias in l-MLLM.
+The objective function is:
 $\mathcal{L}_{\text{D2S}}(\pi_{\text{S}};\pi_{\text{T}})=-\mathbb{E}_{(x,y_{k})\sim\pi_{\text{S}}}\left[\log\pi_{\text{S}}(y_{k}\mid y_{<k},x)\right]-\mathbb{E}_{(x,y_{k})\sim\pi_{\text{T}}}\left[\log\frac{\pi_{\text{T}}(y_{k}\mid y_{<k},x)}{\pi_{\text{S}}(y_{k}\mid y_{<k},x)}\right].$ (7)
 Preference Distillation.
-In this stage, our goal is to distill the preference knowledge from the teacher MLLM to guide the student MLLM towards generating not only accurate but also reasonable responses, which is crucial in reducing hallucinations.
-For the training process, we make effective use of preference data, which comprises meticulously paired positive responses $y^{+}$ and negative responses $y^{-}$ for the identical prompt $x$.
-Our preference distillation strategy is inspired by recent advancements in Direct Preference Optimization (DPO) (Rafailov et al., 2024), which bypasses the need for training a reward model by directly training on an offline preference dataset.
-Our key insight is to treat the teacher MLLM as the reference model during distillation.
-The teacher model plays a crucial role by providing insights into what constitutes “good” and “bad”, thereby establishing a fundamental reference for the student model.
-Specifically, the training objective is to optimize the student model to assign higher probabilities to positive responses than those of the teacher model while assigning lower probabilities to negative responses than those of the teacher model.
-This training process involves two key optimization aspects: First, the student model seeks to align with the teacher model in distinguishing positive from negative responses.
-Second, the student model aims to surpass the teacher model by assigning higher probabilities to positive responses and lower probabilities to negative responses.
-We only train the mixture of experts and the VL adaptor in the student model and employ a Top-$k$ routing strategy to select the experts.
+In this stage, our goal is to distill the preference knowledge from l-MLLM to guide s-MLLM towards generating not only accurate but also reasonable responses, which is crucial in reducing hallucinations.
+For the training process, we effectively use preference data, which comprises meticulously paired positive responses $y^{+}$ and negative responses $y^{-}$ for the identical prompt $x$.
+Our preference distillation strategy is inspired by recent advancements in Preference Optimization (PO) (Rafailov et al., 2024; Ethayarajh et al., 2024), which bypasses the need for training a reward model by directly training on an offline preference dataset.
+Our key insight is to treat l-MLLM as the reference model to provide insights on what constitutes “good” and “bad”, thereby establishing a fundamental reference for s-MLLM.
+Specifically, the training objective is to optimize s-MLLM to assign higher probabilities to positive responses and lower probabilities to negative ones compared to l-MLLM.
+This training process involves two key optimization aspects: First, s-MLLM aims to align with the teacher model in distinguishing positive from negative responses.
+Second, s-MLLM seeks to surpass l-MLLM by assigning higher probabilities to positive responses and lower probabilities to negative responses.
+We only train the experts and adaptor in s-MLLM and employ a Top-$k$ routing strategy to select the experts.
 The trainable parameters are $\theta=\{\omega,\phi_{e}\}$.
-The object function is defined as:
+The objective function is:
 $\mathcal{L}_{\text{PD}}(\pi_{\text{S}};\pi_{\text{T}})=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\text{S}}(y^{+}\mid x)}{\pi_{\text{T}}(y^{+}\mid x)}-\beta\log\frac{\pi_{\text{S}}(y^{-}\mid x)}{\pi_{\text{T}}(y^{-}\mid x)}\right)\right],$ (8)
-where $\pi_{\text{S}}(y^{+}\mid x)$ and $\pi_{\text{S}}(y^{-}\mid x)$ denote the probabilities of positive and negative responses for the student model, respectively, and $\pi_{\text{T}}(y^{+}\mid x)$ and $\pi_{\text{T}}(y^{-}\mid x)$ denote the same for the teacher model.
+where $\pi_{\text{S}}(y^{+}\mid x)$ and $\pi_{\text{S}}(y^{-}\mid x)$ denote the probabilities of positive and negative responses for s-MLLM, and $\pi_{\text{T}}(y^{+}\mid x)$ and $\pi_{\text{T}}(y^{-}\mid x)$ denote the same for l-MLLM.
 4 Experiments
@@
 Implementation Details.
-We employ the well-established ”ViT-MLP-LLM” architecture to demonstrate the effectiveness of LLaVA-MoD.
-Specifically, a pre-trained CLIP-ViT-L/14 is utilized as the vision encoder and a 2-layer MLP is utilized as the adaptor for both the student and the teacher models.
-Both teacher and student models leverage the Qwen-1.5 and Qwen-2 families to construct their base model.
-Specifically, the teacher model employs a 7B parameter configuration, while the student model is constructed with 1.8B and 0.5B parameter sizes.
-The performance of teacher MLLM on common multimodal benchmarks is presented in Table 1.
-Table 1: Performance of teacher MLLM on multimodal benchmarks.
+We employ the “ViT-MLP-LLM” architecture to demonstrate the effectiveness of LLaVA-MoD.
+A pre-trained CLIP-ViT-L/14 is utilized as the vision encoder and a 2-layer MLP is utilized as the adaptor.
+Qwen-1.5/2 with different sizes are utilized as the LLM for l-MLLM and s-MLLM.
+Specifically, l-MLLM is configured with 7B parameters, while s-MLLM is configured with 1.8B and 0.5B parameters.
+The performance of l-MLLM on multimodal benchmarks is presented in Tab.
+1.
+We use the same series of LLMs for distillation, i.e.
+employing Qwen-1.5 7B to distill Qwen-1.5 1.8B and Qwen-1.5 0.5B.
+Each stage uses distinct training configurations.
+The detailed training strategy and hyperparameter are illustrated in Appendix A.1.
+Table 1: Performance of l-MLLM on multimodal benchmarks.
 […]
@@
 Qwen-2-7B CLIP-L 62.5 37.9 73.2 57.2 78.0 71.6 71.1 64.5
-As introduced in Section 3.2, the training process comprises mimic distillation and preference distillation, each with specific settings.
-At initialization, we first freeze the vision encoder and LLM while optimizing the VL adaptor to align image tokens with the word embedding space.
-This stage employs a cross-entropy loss with a batch size of 512 and a learning rate of 1e-4.
-In the mimic distillation stage, the vision encoder remains frozen, while the LLM and VL adaptor are co-optimized to distill general knowledge from the teacher MLLM in a dense-to-dense fashion.
-Then, the FFN in the LLM is first transformed into a sparse architecture with a mixture of FFN experts co-optimized with the VL adaptor to distill specialized knowledge from the teacher MLLM in a dense-to-sparse fashion.
-This stage employs a KL-divergence loss, with cross-entropy loss added for dense-to-sparse distillation.
-The batch size is 256, and the learning rate is adjusted to 2e-5.
-In the preference distillation stage, the model inherits from the mimic distillation.
-The vision encoder remains frozen, and a mixture of FFN experts is co-optimized with the VL adaptor to distill preference knowledge from the teacher MLLM.
-This stage employs a DPO loss to optimize the student’s probability for positive responses to be greater than that of the teacher and the student’s probability for negative responses to be lower than that of the teacher.
-The batch size is 256, and the learning rate is adjusted to 2e-6.
-Throughout all stages, we employ the adam optimizer (Diederik, 2014) and train on 16 NVIDIA A100 GPUs for one epoch each, totaling approximately 960 GPU hours.
-The detailed hyperparameters of each training stage are illustrated in Appendix  11.
 Training Datasets.
 The training data consists of 5M samples from the open-source datasets, with each training stage utilizing a distinct dataset.
-During Initialization, 0.6M general captioning samples from LLaVA-1.5-pretrain dataset (Liu et al., 2023b) are used to bridge the gap between visual and language modalities.
-In mimic distillation, 2.4M general captioning and conversation samples are used to distill general knowledge from the teacher MLLM and 1.4M multi-tasks data, including VQA, documents, science, and OCR, are used to distill specialized knowledge from the teacher MLLM.
-For preference distillation, 0.8M preference data samples are used to transfer preference knowledge from the teacher.
-The detailed dataset of each training stage is illustrated in Appendix 10.
-Evaluation Benchmarks.
+During Initialization, 0.6M general captioning samples are used to bridge the gap between visual and language modalities.
+In mimic distillation, 2.4M general captioning and conversation samples are used to distill general knowledge from l-MLLM, and 1.4M multi-tasks data, including VQA, documents, science, and OCR, are used to distill specialized knowledge from l-MLLM.
+For preference distillation, 80K preference data samples are used to transfer preference knowledge from l-MLLM.
+The detailed dataset of each training stage is illustrated in Appendix A.2.
 […]
-4.2 Main Results
-In this section, we conduct experiments of LLaVA-MoD to highlight its advantages in two aspects: performance and efficiency.
-For performance, we evaluate comprehension-oriented benchmarks (Table 2) and hallucination-oriented benchmarks (Table 3).
-For efficiency, we present a comparison in terms of data samples and model size.
 […]
@@
 InstructBLIP Vicuna-13B 130M 49.5 33.4 63.1 50.7 60.6 - - -
-IDEFICS-80B LLaMA-65B 6.8M 38.4 35.5 59.9 25.9 64.2 48.2 25.2 42.5
 Qwen-VL-Chat Qwen-7B 1450M 57.5 38.9 68.2 61.5 74.4 60.6 56.7 56.7
-SPHINX-Intern2 InternLM2-7B 15M 56.2 49.6 70.4 58.1 63.2 57.9 - -
 Deepseek-VL-7B DLLM-7B 2000M 61.3 49.9 74.0 64.7 73.4 74.1 72.8 67.2
@@
 VILA-3B LLaMA-2.7B 51M 61.5 53.5 69.0 60.4 72.1 63.4 52.7 61.8
-LLaVA-Phi Phi-2-2.7B 1.3M - 35.9 68.4 48.6 66.7 59.8 - -
-TinyLLaVA Phi-2-2.7B 3.2M 61.0 - 70.1 53.5 71.6 68.3 - -
-TinyGPT-V Phi-2-2.7B 24M 33.6 33.4 - - - - - -
 MobileVLM MLLaMA-2.7B 1.3M 59.0 - 61.0 47.5 64.4 59.6 - -
@@
 DeepSeek-VL-1.3B DLLM-1.3B 2000M 59.3 36.8 64.2 58.4 65.3 64.6 61.0 58.5
-LLaVA-MoD-2B Qwen-1.5-1.8B 5M 58.7 39.2 68.0 58.5 66.7 66.3 61.9 59.9
+LLaVA-MoD-2B Qwen-1.5-1.8B 5M 58.7 39.2 68.0 58.5 67.6 66.3 61.9 59.9
 LLaVA-MoD-2B Qwen-2-1.5B 5M 58.8 40.4 69.2 59.9 69.2 68.9 64.4 61.6
@@
 LLaVA-MoD-1B Qwen-2-0.5B 5M 56.6 35.1 61.1 57.1 67.0 58.7 54.1 55.7
-4.2.1 Comprehension-Oriented Benchmarks
-As indicated in Table 2, LLaVA-MoD demonstrates remarkable performance on comprehension-oriented benchmarks.
-It achieves SOTA average results among the models of 2B size and 1B size.
-Specifically, The 2B-sized LLaVA-MoD surpasses Mini-Gemini-2B (Li et al., 2024c) by 8.1%, while using a lower image resolution (336 vs.
+4.2 Main Results
+In this section, we conduct experiments of LLaVA-MoD to highlight its advantages in two aspects: performance and efficiency.
+For performance, we evaluate comprehension-oriented benchmarks (Tab.
+2) and hallucination-oriented benchmarks (Tab.
+3).
+For efficiency, we present a comparison in terms of data samples and model size.
+The performance is obtained under the same series of LLMs, i.e.
+employing Qwen-1.5 7B to distill Qwen-1.5 1.8B and Qwen-1.5 0.5B.
+Comprehension-Oriented Benchmarks.
+As indicated in Tab.
+2, LLaVA-MoD achieves SOTA average results among the models of 1B and 2B size on comprehension-oriented benchmarks.
+The 2B-sized LLaVA-MoD surpasses Mini-Gemini-2B (Li et al., 2024) by 8.1%, while using a lower image resolution (336 vs.
 768).
@@
 Furthermore, LLaVA-MoD-2B matches and even surpasses the performance of large-scale MLLMs.
-Specifically, The 2B-sized LLaVA-MoD surpasses Qwen-VL-Chat-7B (Bai et al., 2023b) by 8.8% and matches the performance of VILA-3B (Lin et al., 2024b) and MiniCPM-V (Yao et al., 2024).
-These results highlight that our approach effectively and efficiently training small-scale MLLMs by distilling the sparse MoE architecture.
+The 2B-sized LLaVA-MoD surpasses Qwen-VL-Chat-7B (Bai et al., 2023b) by 8.8% and matches the performance of VILA-3B (Lin et al., 2024b) and MiniCPM-V (Yao et al., 2024).
+These results highlight that our approach trains small-scale MLLMs effectively by distilling the sparse MoE architecture from large-scale MLLMs.
 Table 3: Comparison with state-of-the-art MLLMs on the hallucination benchmarks.
@@
 The best result is in bold, and the second-best result is underlined.
-Model LLM #Param Object POPE MMHal
-HalBench Bench
+Model LLM #Param Object HalBench POPE MMHal-Bench
 Resp $\downarrow$ Ment $\downarrow$ F1 $\uparrow$ Score $\uparrow$ Hall $\downarrow$
@@
 LLaVA-MoD-2B Qwen-2-1.5B 1.9B 11.2 5.9 87.2 2.91 13.8
-4.2.2 Hallucination-Oriented Benchmarks
-As indicated in Table 3, LLaVA-MoD shows remarkable performance in mitigating hallucination, even beating its teacher model.
-It can be attributed into two aspects: Firstly, by assigning a higher probability for the positive response, preference distillation encourages the student model to focus on providing correct and relevant information.
+Hallucination-Oriented Benchmarks.
+As indicated in Tab.
+3, LLaVA-MoD shows remarkable performance in mitigating hallucination, even beating its teacher model.
+It can be attributed to preference distillation from two aspects: Firstly, by assigning a higher probability for the positive response, preference distillation encourages LLaVA-MoD to focus on providing correct and relevant information.
 Secondly, by assigning a lower probability for the negative response, preference distillation discourages incorrect or unsubstantiated information.
-Using the teacher model as a reference to adjust response probabilities, this optimization enables the student model to handle hallucination problems more accurately and reliably, thereby outperforming the teacher model.
-Moreover, LLaVA-MoD surpasses similar small-sized models (Yao et al., 2024; Li et al., 2024c; He et al., 2024) and even recent RLHF-based models (Sun et al., 2023; Zhou et al., 2024b; Huang et al., 2024; Leng et al., 2024; Zhou et al., 2023; Zhao et al., 2023b).
-For instance, on the Object HalBench, it outperforms RLHF-V (Yu et al., 2024a) by 8.2% in response-level hallucination rate and by 21.3% in mention-level hallucination rate.
-This demonstrates that taking the teacher model as a reference for preference distillation is a challenging task that can encourage the student model to effectively learn to mitigate hallucination.
-4.2.3 Efficiency Comparison
+By using the teacher model as a reference to adjust response probabilities, preference distillation enables LLaVA-MoD to handle hallucination issues more accurately and reliably, thereby surpassing its teacher model.
+Moreover, LLaVA-MoD even surpasses recent RLHF-based models (Sun et al., 2023; Zhou et al., 2024b; Huang et al., 2024; Leng et al., 2024; Zhou et al., 2023).
+On the Object HalBench, it outperforms RLHF-V (Yu et al., 2024a) by 8.2% in response-level hallucination rate and by 21.3% in mention-level hallucination rate.
+This demonstrates that preference distillation is an effective task to mitigate hallucination.
+Efficiency Comparison.
 We compare the efficiency of LLaVA-MoD with recent SOTA MLLMs in terms of training data and trainable parameters.
-As indicated in Table 2, The 2B-sized LLaVA-MoD requires only 5M samples and activates merely a subset of its total 2.2B parameters during training and inference.
-It demonstrates that LLaVA-MoD achieves 8.8% higher performance than Qwen-VL-Chat, using only 23% trainable parameters and 0.3% of training data samples, respectively.
-Moreover, when compared to the small-sized model MiniCPM-V with 2.8B trainable parameters, LLaVA-MoD also achieves superior performance with just 1.6% of the data samples and 2B parameters, underscoring its efficiency.
-In conclusion, LLaVA-MoD enables more efficient utilization of data, parameters, and computational resources.
-This not only results in faster training and inference but also provides a practical solution for deploying high-performance models in resource-constrained environments.
+As indicated in Tab.
+2, The 2B-sized LLaVA-MoD requires only 5M samples and activates merely a subset of its total 2.2B parameters during training and inference.
+It achieves 8.8% higher performance than Qwen-VL-Chat, using only 23% trainable parameters and 0.3% of training data samples.
+Moreover, compared to the small-sized model MiniCPM-V with 2.8B trainable parameters, LLaVA-MoD also achieves superior performance with just 1.6% of the data samples and 2B parameters, underscoring its efficiency.
+LLaVA-MoD enables more efficient utilization of data, parameters, and computational resources.
+This not only results in efficient training and inference but also provides a practical solution for deploying high-performance models in resource-constrained environments.
 4.3 Ablation Study
-In this section, we begin by analyzing the impact of preference distillation.
-Next, we conduct ablation studies on a model trained solely with mimic distillation to investigate the effectiveness of our proposed architecture and training strategy on comprehension-oriented benchmarks.
-We utilize CLIP-ViT-L/14 as the default vision encoder.
-The default large language models (LLMs) are Qwen-1.5-1.8B and Qwen-1.5-7B for the student and teacher models.
+In this section, we thoroughly explore the effect of the distillation strategy and model design.
+Firstly, we investigate the preference distillation on comprehension and hallucination benchmarks.
+Since our focus is on the training strategy and model design, we carry out subsequent ablation experiments specifically on mimic distillation.
+We use CLIP-ViT-L/14 as the default vision encoder, and Qwen-1.5-1.8B and Qwen-1.5-7B as the default student and teacher LLMs respectively.
 4.3.1 Impact of Preference Distillation
-We extensively evaluated preference distillation, examining its impact on both the comprehension ability and hallucination mitigation capabilities of LLaVA-MoD.
-As detailed in Table 4 and Table 5, preference distillation does not yield consistent improvements in comprehension across all benchmarks, its impact on hallucination mitigation was strikingly consistent.
-Across all evaluated benchmarks, preference distillation notably reduced the occurrence of hallucinations, highlighting its efficacy in enhancing the reliability and factual accuracy of LLaVA-MoD.
-This observation aligns with the field of language modeling (Achiam et al., 2023), have consistently indicated that RLHF techniques tend to prioritize the reduction of hallucinations, often at the expense of potential gains in other areas, such as raw comprehension scores.
-Table 4: Performance of preference distillation on comprehension-oriented benchmarks.
+Table 4: The impact of preference distillation on the comprehension-oriented benchmarks.
 […]
@@
 Mimic+Preference 58.7 39.2 68.0 58.5 66.7 66.3 61.9 59.9
-Table 5: Performance of preference distillation on hallucination-oriented benchmarks.
-Distillation Stage Object POPE MMHal
-HalBench Bench
+Table 5: The impact of preference distillation on the hallucination-oriented benchmarks.
+Distillation Object HalBench POPE MMHal-Bench
 Resp $\downarrow$ Ment $\downarrow$ F1 $\uparrow$ Score $\uparrow$ Hall $\downarrow$
@@
 Mimic+Preference 11.4 7.2 87.0 2.76 17.2
+Preference Distillation Mitigates Hallucination.
+We explore the effect of preference distillation on comprehension capability and hallucination mitigation.
+As shown in Tab.
+4 and Tab.
+5, preference distillation remarkably reduces the hallucination of s-MLLM, while it does not yield consistent improvements in comprehension capability.
+This observation aligns with that in LLMs (Achiam et al., 2023), where preference optimization tends to prioritize reducing hallucinations, often resulting in a certain degree of decline in comprehension gains.
+KTO Stabilizes Preference Distillation.
+We explore the training stability of various preference-optimization losses.
+We employ KTO (Ethayarajh et al., 2024) and DPO (Rafailov et al., 2024).
+As shown in Appendix B.2, both methods enhance hallucination mitigation.
+However, KTO has a significant advantage over DPO.
+The result implies that KTO, which indicates whether a sample is good or bad, provides a better signal for s-MLLM to surpass l-MLLM compared to DPO, which determines which one is better by comparing two samples.
 4.3.2 Impact of Training Strategy
-Knowledge Distillation Improves Training Efficiency in MoE.
-This study explores the impact of knowledge distillation (KD) and examines how it facilitates MoE model training compared to supervised fine-tuning (SFT).
-The experiments are conducted using the same sparse configuration of E4T2, where 4 experts are initialized and the top-2 experts are activated.
-The same datasets and training hyperparameters are employed for both KD and SFT.
-Training MoE models is inherently challenging, as it requires effectively leveraging huge data to optimize experts’ routing.
-Our approach demonstrates that KD, compared to traditional SFT, excels in addressing this challenge, achieving superior performance across multiple benchmarks.
-As demonstrated in Table 6, KD significantly outperforms SFT on all the benchmarks.
-Specifically, KD achieves an average improvement of 8.1% over SFT, with particularly pronounced gains in complex multi-task scenarios.
-For instance, on the MMB and MME benchmarks, KD surpasses SFT by 8.2% and 10.0%, respectively.
-These results highlight KD’s ability to effectively leverage multiple experts and extract comprehensive knowledge from large-scale MLLMs.
-By distilling essential knowledge, KD not only maximizes data utilization but also achieves better generalization ability of MoE compared to SFT.
-Table 6: Comparison between the KD and SFT. The architecture of the student model is configured as E4T2.
+KD Facilitates MoE Training.
+We explore the effect of knowledge distillation (KD) in MoE training by comparing it with supervised fine-tuning (SFT).
+The ablation experiments are carried out using the same sparse configuration of E4T2, where four experts are initialized and the top two experts are activated.
+As shown in Tab.
+6, KD surpasses SFT on all benchmarks, achieving an 8.1% average gain over SFT.
+Additionally, it achieves notable gains in complex multi-task scenarios, such as a +8.2% gain on MMB and a +10.0% gain on MME.
+These results indicate that KD provides a better optimization signal for effectively training the MoE architecture.
+Table 6: Comparison between KD and SFT. The MoE architecture is E4T2.
 […]
@@
 KD 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
-Focusing on Response Improves Distillation Generalization.
-This study explores the impact of the distillation target.
-The default approach distills response, we experiment with distilling both response and instruction.
-Both settings utilize the same datasets and training hyperparameters.
-Table 7 shows that distilling only the response achieves superior performance, surpassing jointly distilling response and instruction by 3.1% in terms of average performance.
-This suggests that focusing on response is more effective for knowledge distillation in auto-regressive modeling.
-A possible explanation lies in that mimicking instructions could lead to overfitting the specific instruction patterns of training data, potentially reducing the model’s adaptability to unseen instructions.
-Mimicking the teacher model’s response proves to be effective in improving the student’s auto-regressive modeling generalization capabilities.
-Table 7: Comparison between different distillation tokens.
-The response indicates distilling solely on the generated response tokens, while the response+instruction incorporates additional instruction tokens in the distillation process.
+General-to-Specialized KD Boosts Performance.
+We explore the effect of the general-to-specialized approach in mimic distillation by comparing the proposed D2D+D2S with D2S.
+The D2D+D2S involves a two-phase process.
+Firstly, it distills a dense s-MLLM using general data.
+Subsequently, it transforms the s-MLLM from dense to sparse and further distills it using multi-task data.
+The D2S directly distills the sparse s-MLLM using a combined of general and multi-task dataset.
+Tab.
+7 shows that D2D+D2S surpasses D2S with an average gain of 4.0%.
+The result indicates that D2D+D2S enables effective knowledge transfer, striking a balance between general and specialized competencies.
+Additionally, D2D+D2S is more computationally efficient, consuming only 62.5% of the GPU hours required by D2S.
+Table 7: Comparison between different distillation strategies in mimic distillation. #GPU indicates the training GPU hours.
+[…]
+D2S 1536 57.9 36.5 66.3 57.5 65.7 61.4 60.0 57.9
+D2D+D2S 960 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
+Focus on Response Distillation Improves Generalization.
+We explores the effect of the distillation target by comparing response distillation with response+instruction distillation.
+Tab.
+8 shows that response distillation surpasses response+instruction distillation with an average gain of 3.1%.
+The result indicates that focusing on response is more effective for knowledge distillation in auto-regressive modeling.
+A possible reason for this is that mimicking additional instructions can lead to overfitting the specific instruction patterns in the training data, potentially reducing the generalization to unseen instructions.
+Table 8: Comparison between different distilled tokens. Response indicates distilling solely on the response tokens. Response+Instruction incorporates additional distillation of instruction tokens.
 […]
@@
 4.3.3 Impact of Model Architecture
-Sparse MoE Enables Effective Knowledge Distillation.
-This study explores the effect of sparse architecture in distillation by comparing its performance to dense architecture.
-We adopt the E4T1 configuration as the sparse architecture to ensure the same activated parameters as the dense architecture.
-The same datasets and training hyperparameters are used for sparse and dense architectures.
-[…]
-Table 8: Comparison between sparse and dense architecture within the distillation. The sparse architecture is configured as E4T1.
+Sparse Architecture Facilitates Knowledge Transfer.
+We explore the effect of sparse architecture in distillation by comparing it with dense architecture.
+The configuration is E4T1 where four experts are initialized and the top-1 expert is activated to ensure that the activated parameters are the same as those of the dense architecture.
+Tab.
+9 shows that sparse architecture surpasses dense architecture with an average gain of 3.7%, The superior performance is prominent in complex multi-task benchmarks, such as MME and MMB.
+The result indicates that sparse architecture leverages MoE to effectively transfer diverse knowledge from l-MLLM while maintaining computational efficiency.
+Table 9: Comparison between sparse and dense architecture within the distillation.
 […]
@@
 Teacher Capacity Matters in Knowledge Distillation.
-This study explores the impact of the teacher in the distillation process.
-Specifically, we utilize CLIP-ViT-L/14 as the vision encoder.
-For the LLM, Qwen-1.5-7B is the strong teacher while Qwen-1.5-4B is the weaker teacher.
-All experiments are conducted under the same datasets and training hyperparameters to ensure a fair comparison.
-As shown in Table 9, a suitable teacher enhances performance across various benchmarks.
-For instance, the Qwen-1.5-1.8B student model achieves an average performance of 62.3% with the weaker Qwen-1.5-4B teacher.
-This performance jumps to 63.6% when utilizing the stronger Qwen-1.5-7B teacher.
-Interestingly, the smaller Qwen-1.5-0.5B student model shows marginal improvement with either the Qwen-1.5-4B or Qwen-1.5-7B teacher.
-This suggests that an excessively large capacity gap between the teacher and student models can hinder effective knowledge transfer.
-Utilizing a ”middle teacher” with an intermediate capacity could bridge this gap, facilitating smoother knowledge transfer and boosting the student model’s learning efficiency.
-Table 9: Comparison between the strong and weak teachers within the distillation.
-We set the strong teacher as the LLM is Qwen-1.5-7B and the weak teacher as the LLM is Qwen-1.5-4B.
-The model configuration is E4T2.
+We explore the effect of the teacher capacity in the distillation.
+We employ Qwen-1.5-7B as the strong teacher and Qwen-1.5-4B as the weaker one.
+The MoE configuration is set as E4T2.
+Tab.
+10 shows that a well-suited teacher can boost performance.
+For Qwen-1.5-1.8B, using a 7B-sized teacher yields an average gain of 2.2% compared to a 4B-sized teacher.
+However, when the student and teacher have a large capacity disparity, it can hinder effective knowledge transfer.
+For Qwen-1.5-0.5B, using a 7B-sized teacher yields a marginal average gain compared to a 4B-sized teacher.
+Utilizing a “middle teacher” with an intermediate capacity can bridge the gap, facilitating smooth knowledge transfer.
+Table 10: Comparison between the strong and weak teachers. We set the strong teacher as the LLM is Qwen-1.5-7B and the weak teacher as the LLM is Qwen-1.5-4B.
 […]
@@
 Extensive experiments show that LLaVA-MoD outperforms existing models with low activation parameters and computational costs.
-Notably, it outperforms Qwen-VL-Chat-7B by 8.8% with only 2 billion activation parameters, 0.3% training data, and 23% trainable parameters, highlighting its effectiveness in distilling knowledge and driving more efficient multimodal language model development.
-6 Limitations
-LLaVA-MoD requires that both the teacher and student models belong to the same family of large language models (LLMs) to ensure consistency in the vocabulary space.
-Future research could explore distillation techniques involving heterogeneous model families.
-Additionally, the need to load both student and teacher models results in significant memory usage.
-To facilitate more efficient distillation, a feasible solution would be to pre-extract the outputs of the teacher model, allowing only the student model to be loaded during training.
-Appendix A Appendix
-A.1 Dataset
-We report the detailed training dataset settings of LLaVA-MoD in Table 10.
-Table 10: Training dataset. #Sample means the training samples.
-Stage Task Dataset
-Initialization Captioning LLaVA-1.5-Pretrain (Liu et al., 2023b)
-Dense-to-Dense Distillation Captioning ALLaVA-Caption-4V (Chen et al., 2024), ShareGPT4V-PT (Chen et al., 2023a)
-Conversation MIMIC-IT (Li et al., 2023a), LVIS (Wang et al., 2023a), LRV (Liu et al., 2023a), SViT (Zhao et al., 2023a)
-Dense-to-Sparse Distillation Captioning ShareGPT4V-100K (Chen et al., 2023a), TextCaps (Sidorov et al., 2020)
-Conversation LLaVA-1.5-Instruct (Liu et al., 2023b)
-General QA GQA (Hudson & Manning, 2019), VQAv2 (Goyal et al., 2017), OKVQA (Marino et al., 2019)
-Grounding Visual Genome (Krishna et al., 2017), RefCOCO (Yu et al., 2016; Mao et al., 2016)
-Science AI2D (Kembhavi et al., 2016), ScienceQA (Lu et al., 2022a)
-Chart & Doc DVQA (Kafle et al., 2018), ChartQA (Masry et al., 2022), DocQA (Clark & Gardner, 2018)
-OCR OCRVQA (Mishra et al., 2019), SynthDoG-EN (Kim et al., 2022)
-Knowledge A-OKVQA (Schwenk et al., 2022), GeoQA+ (Cao & Xiao, 2022)
-Preference Distillation Preference RLAIF-V (Yu et al., 2024b)
-A.2 Hyperparameters
-We report the detailed training hyperparameters of LLaVA-MoD in Table 11.
-Table 11: Training hyperparameters
-Configuration Stage-1 Stage-2 Stage-3
+Notably, it outperforms Qwen-VL-Chat-7B by 8.8% with only 2 billion activation parameters, 0.3% training data, and 23% trainable parameters, highlighting its effectiveness in distilling knowledge and driving more efficient MLLM development.
+Appendix A Implementation Details
+A.1 Training Strategy and Hyperparameters
+We first freeze the vision encoder and LLM while optimizing the VL adaptor to align image tokens with the word embedding space during initialization.
+This stage employs a cross-entropy loss with a batch size of 512 and a learning rate of 1e-4.
+In mimic distillation, the vision encoder remains frozen, while the LLM and VL adaptor are co-optimized to distill general knowledge from l-MLLM in dense-to-dense distillation.
+Then, the FFN in the LLM is first transformed into a sparse architecture with a mixture of FFNs co-optimized with the VL adaptor to distill specialized knowledge from l-MLLM in dense-to-sparse distillation.
+This stage employs a KL-divergence loss, with cross-entropy loss added for dense-to-sparse distillation.
+The batch size is 256, and the learning rate is adjusted to 2e-5.
+In preference distillation, the model inherits from the mimic distillation.
+The vision encoder remains frozen, and a mixture of FFN experts is co-optimized with the VL adaptor to distill preference knowledge from the teacher MLLM.
+This stage employs a KTO (Ethayarajh et al., 2024) loss to optimize the probability of s-MLLM for positive responses to be greater than that of l-MLLM and the probability of s-MLLM for negative responses to be lower than that of l-MLLM.
+The batch size is 256, and the learning rate is adjusted to 2e-6.
+Throughout all stages, we employ the adam optimizer (Diederik, 2014) and train on 16 NVIDIA A100 GPUs for one epoch each, totaling approximately 960 GPU hours.
+The detailed training hyperparameters are shown in Tab.
+11.
+Table 11: Training hyperparameters of each stage.
+Configuration Initialization Mimic Distillation Preference Distillation
 LLM
@@
 Optimizer AdamW
-Optimizer hyperparameter $\beta_{1}=0.9,\beta_{2}=0.98,eps=1e^{-6}$
+Optimizer hyperparameter $\beta_{1}=0.9,\beta_{2}=0.98$
 Learning rate 1e-4 2e-5 2e-5
@@
 Model parallelism Zero2 Zero2 offload Zero2 offload
+Table 12: Training dataset of each stage. #Sample means the training samples
+Stage Task Dataset
+Initialization Captioning LLaVA-1.5-Pretrain (Liu et al., 2023b)
+Dense-to-Dense Distillation Captioning ALLaVA-Caption-4V (Chen et al., 2024), ShareGPT4V-PT (Chen et al., 2023a)
+Conversation MIMIC-IT (Li et al., 2023a), LVIS (Wang et al., 2023), LRV (Liu et al., 2023a), SViT (Zhao et al., 2023)
+Dense-to-Sparse Distillation Captioning ShareGPT4V-100K (Chen et al., 2023a), TextCaps (Sidorov et al., 2020)
+Conversation LLaVA-1.5-Instruct (Liu et al., 2023b)
+General QA GQA (Hudson & Manning, 2019), VQAv2 (Goyal et al., 2017), OKVQA (Marino et al., 2019)
+Grounding Visual Genome (Krishna et al., 2017), RefCOCO (Yu et al., 2016; Mao et al., 2016)
+Science AI2D (Kembhavi et al., 2016), ScienceQA (Lu et al., 2022a)
+Chart & Doc DVQA (Kafle et al., 2018), ChartQA (Masry et al., 2022), DocQA (Clark & Gardner, 2018)
+OCR OCRVQA (Mishra et al., 2019), SynthDoG-EN (Kim et al., 2022)
+Knowledge A-OKVQA (Schwenk et al., 2022), GeoQA+ (Cao & Xiao, 2022)
+Preference Distillation Preference RLAIF-V (Yu et al., 2024b)
+A.2 Training Datasets
+We curate a comprehensive open-source dataset of 5 million samples encompassing both general and expert tasks.
+In the Initialization stage, we utilize 0.6M general captioning samples from LLaVA-1.5-pretrain dataset (Liu et al., 2023b) to bridge the gap between visual and language modalities.
+In the mimic distillation stage, we utilize 2.4M general captioning and conversation samples to distill general knowledge from the teacher MLLM and 1.4M multi-tasks data samples, including VQA, documents, science, and OCR to distill specialized knowledge from the teacher MLLM.
+In the preference distillation stage, we utilize 80K preference data samples to transfer preference knowledge from the teacher MLLM.
+The detailed dataset of each training stage is illustrated in Tab.
+12.
+Appendix B Additional Experiments
+B.1 The Effect of Expert and Router Capacity
+We explore the effects of expert and router capacity on the results.
+Specifically, we utilize CLIP-ViT-L/14 as the vision encoder and the Qwen-1.5-1.8B as the LLM.
+We experiment with experts of 4 and 8, and router strategies of Top-1 and Top-2.
+All experiments are conducted under the same datasets and training hyperparameters to ensure a fair comparison.
+Tab.
+14 shows that scaling experts from 4 to 8 doesn’t improve performance.
+Tab.
+14 shows that scaling the router strategy from Top-1 to Top-2 improves performance.
+This indicates that increasing the capacity of the MoE block could introduce training difficulty.
+As the capacity grows, the model might struggle more with convergence and optimization, potentially making it harder to effectively learn from the given data.
+Thus, a careful balance must be achieved between capacity and training efficiency to ensure optimal performance.
+Table 13: Ablations with the number of experts. We set the gating strategy as Top-2. The number of experts is set to 4 and 8.
+[…]
+4 59.3 68.4 58.7 68.4 65.1 64.0
+8 58.3 67.4 58.4 69.1 64.6 63.6
+Table 14: Ablations with the router strategy. We set the number of experts as 4. The Top-K strategy of routers is set to Top-1 and Top-2.
+[…]
+1 58.2 67.2 58.3 66.1 64.7 62.6
+2 59.3 68.4 58.7 68.4 65.1 64.0
+B.2 The Effect of Preference-Optimization Loss in Preference Distillation
+We explore the training stability of various preference-optimization losses.
+We employ KTO (Ethayarajh et al., 2024) and DPO (Rafailov et al., 2024).
+As shown in Tab.
+15, while both methods have an impact on reducing hallucination, KTO is significantly superior to DPO.
+This is because KTO, which indicates whether a sample is good or bad, offers a better signal for the s-MLLM to surpass the l-MLLM compared to DPO, which determines relative quality by comparing two samples.
+Table 15: Comparison between different preference-optimization loss on comprehension-oriented benchmarks.
+[…]
+DPO 58.5 37.3 68.2 58.6 67.8 65.4 61.5 59.6
+KTO 58.7 39.2 68.0 58.5 67.6 66.3 61.9 59.9
+Table 16: Comparison between different preference-optimization loss on hallucination-oriented benchmarks.
+PO Loss Object HalBench POPE MMHal-Bench
+Resp $\downarrow$ Ment $\downarrow$ F1 $\uparrow$ Score $\uparrow$ Hall $\downarrow$
+DPO 20.0 10.5 86.8 2.68 18.9
+KTO 11.4 7.2 87.0 2.76 17.2
+Appendix C Limitations and Future Works
+LLaVA-MoD requires that both s-MLLM and l-MLLM belong to the same series of LLMs to ensure consistency in the vocabulary space.
+Future research can explore distillation techniques involving heterogeneous model families.
+Moreover, the requirement to load both s-MLLM and l-MLLM leads to substantial memory consumption.
+To enable efficient distillation, a possible solution could be to pre-extract the logits of the l-MLLM.
+This would allow only the s-MLLM to be loaded during training, reducing memory requirements and potentially speeding up the training process.